### v1

# Universality of Estimator for High-Dimensional Linear Models with Block Dependency

Abstract.
We study the universality property of estimators for high-dimensional linear models, which exhibits a distribution of estimators is independent of whether covariates follows a Gaussian distribution. Recent high-dimensional statistics require covariates to strictly follow a Gaussian distribution to reveal precise properties of estimators. To relax the Gaussianity requirement, the existing literature has studied the conditions that allow estimators to achieve universality. In particular, independence of each element of the high-dimensional covariates plays an important role. In this study, we focus on high-dimensional linear models and covariates with block dependencies, where elements of covariates only within a block can be dependent, then show that estimators for the model maintain the universality. Specifically, we prove that a distribution of estimators with Gaussian covariates is approximated by an estimator with non-Gaussian covariates with same moments, with the setup of block dependence. To establish the result, we develop a generalized Lindeberg principle to handle block dependencies and derive new error bounds for correlated elements of covariates. We also apply our result of the universality to a distribution of robust estimators.
1. Introduction
We consider an estimator for a linear model by minimizing an empirical risk. Let $p\in\mathbb{N}$ be dimension and $\mathcal{Y}$ is a space of output variables. Let $n\in\mathbb{N}$ be a number of observations, and $(X_{1},Y_{1}),...,(X_{n},Y_{n})$ be a set of independently and identically distributed $\mathbb{R}^{p}\times\mathcal{Y}$-valued random observations generated from a statistical model.
$Y_{i}=X_{i}^{\top}\theta_{0}+\xi_{i},~{}i=1,...,n,$ (1)
where $\theta_{0}\in\mathbb{R}^{p}$ is a unknown parameter and $\xi_{i}$ is an independent noise variable. With a convex loss function $\ell_{0}:\mathbb{R}\to\mathbb{R}$ and a convex regularizer $R_{0}:\mathbb{R}^{p}\to\mathbb{R}$, we consider an estimator $\widehat{\theta}_{\mathbf{X}}$ through the empirical risk minimization problem:
$\widehat{\theta}_{\mathbf{X}}\in\operatornamewithlimits{argmin}_{\theta\in\mathbb{R}^{p}}\left\{\frac{1}{n}\sum_{i=1}^{n}\ell_{0}(Y_{i}-X_{i}^{\top}\theta)+R_{0}(\theta)\right\}.$ (2)
In this study, we show the universality of the distribution of this estimator $\widehat{\theta}_{\mathbf{X}}$, that is, a distribution of the estimator does not depend on whether the covariate $X_{i}$ follow a Gaussian distribution. Specifically, we will show that the universality holds even in the presence of weak dependencies between elements of the random covariate $X_{i}$.
The Gaussianity of covariates is a particularly important property in the analysis of high-dimensional statistics, which has been rapidly developing in recent years. Here, the term high-dimensional refers to the so-called proportional high-dimensional regime, where both the number of parameters $p$ in the statistical model and the amount of data $n$ used for estimation diverge while maintaining a certain ratio. In this setting, several advanced frameworks allow the precise determination of the risk and asymptotic distribution of estimators. These tools include not only the random matrix theory (Marchenko and Pastur, 1967; Dobriban and Wager, 2018; Hastie et al., 2022) but also frameworks such as the convex Gaussian minimax theorem (Stojnic, 2013; Thrampoulidis et al., 2015, 2018; Miolane and Montanari, 2021; Salehi et al., 2019; Montanari et al., 2019; Loureiro et al., 2021; Liang and Sur, 2022; Celentano et al., 2023; Koehler et al., 2021; Tsuda and Imaizumi, 2023), the approximate massage passing (Donoho et al., 2009; Bayati and Montanari, 2011a, b; Rangan, 2011; Javanmard and Montanari, 2013; Donoho and Montanari, 2016; Sur and Candès, 2019; Barbier et al., 2019; Mondelli and Venkataramanan, 2021; Zhao et al., 2022; Feng et al., 2022; Sawaya et al., 2023), the leave-one-out (El Karoui et al., 2013; Bean et al., 2013; El Karoui, 2018; Sur et al., 2019; Yadlowsky et al., 2021), and others (Mézard et al., 1987; Mezard and Montanari, 2009; Bellec, 2022; Sawaya et al., 2024), which have been applied to analyze the properties of statistical models across a wide range of domains. Although these developments are remarkable, many of these frameworks heavily rely on the strict Gaussianity of covariates, which poses a practical limitation.
Relaxing the Gaussianity of covariates of these high-dimensional statistics is crucial to enhance the practicality of high-dimensional methods. The property that the behavior of statistics or estimators does not rely on Gaussianity is known as universality. Specifically, the universality means that properties of statistics under non-Gaussian covariates asymptotically matches that under Gaussian covariates with the same moments. Several studies (Korada and Montanari, 2011; Montanari and Nguyen, 2017; Hu and Lu, 2022; Goldt et al., 2022; Montanari and Saeed, 2022; Dudeja et al., 2022; Lahiry and Sur, 2024) have shown that, in a risk minimization problem, a value of the minimized risk using high-dimensional models achieves universality. The universality of approximate massage passing algorithms is also extensively studied using a rotationally invariant matrix (Fan, 2022; Dudeja et al., 2023; Wang et al., 2024). The universality has been studied not only in terms of risks but also in detailed properties of estimators required for statistical inference. Han and Shen (2023) shows that the estimators of high-dimensional parameters obtained by risk minimization achieve the universality. Han (2024) demonstrates that an estimator given by a first-order algorithm achieves the universality. Furthermore, the universality has been shown to hold in more specific high-dimensional models.
A challenge in the universality of estimators lies in the requirement on properties of non-Gaussian covariates, particularly the independence between elements of covariates. In Han and Shen (2023) and Han (2024), it is stated that, for the universality of estimators, each element of high-dimensional covariates used for the estimator should be independent. This condition not only fails to hold in real high-dimensional covariates but also poses an obstacle when applying universality results to various statistical models. Related to this issue, Lahiry and Sur (2024) demonstrates that the universality of risks holds under the assumption that elements of $p$-dimensional covariates have weak dependence, specifically block dependence. However, this relaxation of the independence of elements for the universality of estimators remains an ongoing challenge.
In this study, we show that the universality holds for estimators obtained by risk minimization, when elements of covariates exhibit dependencies. Specifically, we consider block dependency of elements of covariates, where each element of covariate vectors depends on at most $d$ other elements. We show that when the dependency coefficient $d$ has a certain polynomial order in $n$ or $p$, a distribution of estimators does not rely on the Gaussianity of covariates, in the limit where $n$ and $p$ diverge while maintaining a proportional relationship. This research follows the universality framework of estimators established by Han and Shen (2023), which allows the dependency considered in Lahiry and Sur (2024). As an application of the result, we demonstrate that we derive the universality of a distribution of robust estimators obtained by minimizing the absolute loss and the Huber loss.
As technical aspects, we develop two proof techniques to demonstrate the universality results above. The first is a generalized Lindeberg principle tailored to address the block dependence. In our approach, we measure the discrepancy of two design matrices by focusing on a cell of dependent elements of covariates. We then show that this error is bounded by a cubic polynomial of the number of correlated elements under block dependence, which is critical to controlling the order of block dependence. The second technique involves deriving an upper bound on the number of dependent elements under block dependence. We generate a refined partition based on the original partition of elements, then we derive a tighter upper bound from the new partition. These two techniques allow for a precise evaluation of errors compared to the previous studies, thereby enabling the universality under block dependence with the larger order.
We mention the differences between our research and related studies. Lahiry and Sur (2024) is the closest to our work, which develops the notion of block dependence and derives the universality of risks under block dependence. While Lahiry and Sur (2024) utilizes Stein’s lemma to show the universality of risks, we developed a proof using the Lindeberg approach which enables us to show the universality of estimators. By this approach, we show the universality of the distribution of specific estimators, such as the robust estimator.
1.1. Notation
[…]
2. Setup with Block Sparsity
2.1. Setup: Estimator with Convex Risk Minimization
We consider an estimator with empirical loss minimization under general convex loss and convex regularization. Recall that the random $n$ observations $(X_{1},Y_{1}),...,(X_{n},Y_{n})$ follow a statistical model. Also, we assume that the $\mathbb{R}^{p}$-valued random vectors $X_{1},...,X_{n}$ follows $\mathbb{E}[X_{i}]=0$ and $\mathbb{E}[X_{i}X_{i}^{\top}]=\Sigma/n$ with a positive definite matrix $\Sigma\in\mathbb{R}^{p\times p}$. Also, recall that the estimator $\widehat{\theta}_{\mathbf{X}}$ defined in (2) with the convex loss function $\ell_{0}$ and the regularizer $f_{0}$. We further define design matrices $\mathbf{X}=(X_{1},...,X_{n})^{\top}\in\mathbb{R}^{n\times p}$ of the covariates.
For simplicity, we rewrite the minimization problem (2) and consider an equivalent problem. We define another minimizer $\widehat{w}_{\mathbf{X}}$ by the redefined version of the minimization problem as follows:
$\widehat{w}_{\mathbf{X}}\in\operatornamewithlimits{argmin}_{w\in\mathbb{R}^{p}}H_{\ell}(w,\mathbf{X}):=\operatornamewithlimits{argmin}_{w\in\mathbb{R}^{p}}\left\{\frac{1}{n}\sum_{i=1}^{n}\ell_{i}(X_{i}^{\top}w)+R(w)\right\}.$ (3)
The minimizer has an explicit relation $\widehat{w}_{\mathbf{X}}=\widehat{\theta}_{\mathbf{X}}-\theta_{0}$ and $\ell_{i}(X_{i}^{\top}w)$ and $R(w)$ in (3) correspond to $\ell_{0}(Y_{i}-X_{i}^{\top}\theta)=\ell_{i}(X_{i}^{\top}w)$ and $R_{0}(\theta_{0}+w)$ in (2), respectively.
2.2. Basic Assumption
We introduce assumptions on the dimensionality $p$ and the loss function $\ell_{i}$. These assumptions are common in high-dimensional statistics and allow for a wide class of estimators.
Assumption 1 (Proportionally high-dimension).
There exists $\tau\in(0,1)$ such that $\tau\leq p/n\leq 1/\tau$.
This assumption restricts the high-dimensional setting we consider to the proportionally high-dimensional regime. Since $\tau$ ranges in an open interval $(0,1)$, we can also consider the setting as $p<n$.
Assumption 2 (Loss function).
There exist non-negative real numbers $q_{\ell}$ ($\ell=0,1,2,3$), constants $\bar{\rho}\in(0,1]$, and ${L_{\ell_{i}}\geq 1:i\in[m]}$, along with two measurable functions $\mathcal{D}_{\ell}(\rho)$ and $\mathcal{M}_{\ell}(\rho):(0,\bar{\rho})\rightarrow\mathbb{R}_{\geq 0}$, where $\mathcal{M}_{\ell}(\rho)\leq 1\leq\mathcal{D}_{\ell}(\rho)$, that satisfy the following conditions:
  1. (i)
    The functions $\ell_{i}$ grow at most polynomially, meaning for $i\in[m]$,
    $\sup_{x\in\mathbb{R}}\frac{|\ell_{i}(x)|}{1+|x|^{q_{0}}}\leq L_{\ell_{i}}.$
  2. (ii)
    There exist smooth approximations $\{\ell_{i;\rho}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}\}_{\rho\in(0,\bar{\rho})}$ of $\ell_{i}$ such that (i) $\ell_{i;\rho}\in C^{3}(\mathbb{R})$, (ii) $\max_{i\in[m]}\|\ell_{i;\rho}-\ell_{i}\|_{\infty}\leq\mathcal{M}_{\ell}(\rho)$, and (iii) the derivatives of ${\ell_{i;\rho}}$ adhere to the following self-bounding property:
    $\max_{\ell=1,2,3}\max_{i\in[m]}\sup_{x\in\mathbb{R}}\frac{|\partial\ell_{i;\rho}(x)|}{1+|\ell_{i;\rho}(x)|^{q_{\ell}}}\leq\mathcal{D}_{\ell}(\rho).$
In addition, we define $\mathbf{q}:=\max\{3q_{1},q_{1}+q_{2},q_{3}\},\bar{\mathbf{q}}_{0}:=2\lceil{q_{0}\mathbf{q}/2\rceil}$, and $Av(\{L_{\ell_{i}}^{\mathbf{q}}\}):=(1/n)\sum_{i=1}^{n}L_{\ell_{i}}^{\mathbf{q}}$. Moreover, $q$ refers to the set $\{q_{0},q_{1},q_{2},q_{3},\mathbf{q},\bar{\mathbf{q}}_{0}\}$.
Assumption 2 is exactly the same as an assumption on loss functions in Han and Shen (2023). $\mathcal{M}_{\ell}(\rho)$ and $\mathcal{D}_{\ell}(\rho)$ prescribe the smooth approximation of a loss function. In particular, when a loss function is in $C^{3}(\mathbb{R})$, $\mathcal{M}_{\ell}(\rho)$ is equal to 00. Assumption 2-(i) prevents loss functions from growing too fast as $|x|$ goes to infinity. In many cases, $L_{\ell_{i}}$ is characterized by a noise term of a statistical model. Assumption 2-(ii) restricts on the smoothness of loss functions. It allows loss functions to have a smooth approximation with respect to the supremum norm that increases at most polynomially as in Assumption 2-(ii) . We provide several common examples and their corresponding parameters $q$:
Example 1 (Linear regression with convex losses).
We consider the linear regression model $Y_{i}=X_{i}^{\top}\theta_{0}+\xi_{i}$ with a true parameter $\theta_{0}\in\mathbb{R}^{p}$ and an independent noise variable $\xi_{i}$. With this model, we consider the following loss functions:
  • Squared loss: We set $\ell_{0}(y,z)=(y-z)^{2}$ and $\ell_{i}(z)=(z-\xi_{i})^{2}$, then we obtain $q_{0}=2,\mathbf{q}=3/2$, and $\bar{\mathbf{q}}_{0}=4$. Furthermore, for any $\rho>0$, we have $\mathcal{M}_{\ell}(\rho)=0$ and $\mathcal{D}_{\ell}(\rho)\leq C$ with some constant $C>1$.
  • Absolute loss: We set $\ell_{0}(y,z)=|y-z|$ and $\ell_{i}(z)=|z-\xi_{i}|$, then we obtain $q_{0}=1,\mathbf{q}=0$, and $\bar{\mathbf{q}}_{0}=0$. Furthermore, for any $\rho>0$, we have $\mathcal{M}_{\ell}(\rho)\leq\rho$ and $\mathcal{D}_{\ell}(\rho)\leq C\rho^{-2}$ with some constant $C>1$.
  • Huber loss: We set $\ell_{0}(y,z)=(y-z)^{2}\textbf{1}\{|y-z|\leq\eta\}+(\eta|y-z|-\eta^{2}/2)\textbf{1}\{|y-z|>\eta\}$ with $\eta>0$, then we obtain $q_{0}=1,\mathbf{q}=0$, and $\bar{\mathbf{q}}_{0}=3$. Furthermore, for any $\rho>0$, we have $\mathcal{M}_{\ell}(\rho)\leq\rho$ and $\mathcal{D}_{\ell}(\rho)\leq C\rho^{-2}$ with some constant $C>1$.
For the convex regularizer $R$, we introduce the following continuity measure: for $L,\delta>0$, we define
$\mathcal{M}_{R}(L,\delta):=\sup_{w,w^{\prime}\in[-L,L]^{p}:\|w-w^{\prime}\|_{\infty}\leq\delta}|R(w)-R(w^{\prime})|.$ (4)
This term measures the supremum of discrepancy in the sphere with radius $\delta$ in terms of the uniform norm.
We introduce assumptions about the distribution of the covariate $X_{i}$. The assumption is for moments of the distribution of $X_{i}$ and the largest eigenvalue of the covariance matrix $\Sigma/n$ of $X_{i}$.
Assumption 3 (Distribution of covariates).
Let $q^{*}$ equal to $2^{(\bar{\mathbf{q}}_{0}+4)/2}$ and assume there exists $X_{0}$ such that $X_{i}=X^{0}_{i}/\sqrt{n}$ and the following inequality hold:
$M:=\max$$\left\{\max_{1\leq j\leq p}\|X^{0}_{i,j}\|_{q^{*}},\lambda_{\max}(\Sigma)\right\}<\infty.$ (5)
In the universality studies, controlling the moments and the boundedness of the largest eigenvalue are both common assumptions (Montanari and Saeed, 2022; Han and Shen, 2023; Lahiry and Sur, 2024).
2.3. Dependence structure
We introduce the notion of block dependence to describe the dependencies among elements of a $p$-dimensional random vector $X_{i}$. In preparation, we define a $k$-partition $\{B_{j}\}_{j=1}^{k}$ of a set $[p]$, which is a sequence of $k$ cells $B_{1},...,B_{k}\subset[p]$ that are disjoint to each other and satisfy $\cup_{j=1}^{k}B_{j}=[p]$.
Definition 1 (Block dependence).
A distribution of a $p$-dimensional random vector $X$ is called $d$-block dependent with a dependence parameter $d$, if there exists $k\leq p$ and a $k$-partition $\{B_{j}\}_{j=1}^{k}$ of $[p]$ such that $\max_{j=1,...,k}|B_{j}|\leq d$ and $X_{\ell}\perp\!\!\!\perp X_{k}$ holds for $\ell\in B_{i},k\in B_{j},i\neq j$.
The above dependent structure assigns each element of $X$ to a specific cell $B_{j}$ and restricts the maximum size of $B_{j}$ to $d$. For any two random elements of $X_{i}$ in the same cell $B_{j}$, the structure allows them to be dependent while two random variables that come from different cells must be independent. Note that the dependent elements need not be adjacent to each other.
In the following sections, we consider the situation where the observations $X_{1},...,X_{n}$ follows the distribution which satisfies the $d$-block dependence. Note that the observations are independently and identically distributed, hence the partition $\{B_{j}\}_{j=1}^{k}$ is independent of $i=1,...,n$.
Remark 1 (In related studies).
In several high-dimensional studies, block dependence and related concepts have been utilized. Han and Shen (2023) and Han (2024), which deal with the universality of estimators, assumes the independence of all elements, which can be considered as block dependence with $d=1$. The block dependence with $d\geq 2$ was introduced in Lahiry and Sur (2024), demonstrating universality under $d=o((n/\log n)^{1/5})$. In the field of high-dimensional central limit theorems, Chang et al. (2024) introduces $m$-dependence, a related notion of the block dependence.
3. Main Result of Universality
3.1. Gaussian Analogy
We present the main result, the universality theorem, which demonstrates that the properties of the estimator $\widehat{\theta}_{\mathbf{X}}$ are asymptotically equivalent to those in the case of Gaussian covariates. As a preparation, we define the corresponding Gaussian covariates to $\mathbf{X}$, as well as the risk and estimator for it.
Let $G_{1},...,G_{n}$ be a set of i.i.d. vectors from $N(0_{p},\Sigma/n)$, and its design matrix $\mathbf{G}=(G_{1},...,G_{n})^{\top}\in\mathbb{R}^{n\times p}$. Using the Gaussian vectors, we define Gaussian analogies of the empirical risk and the estimator as
$\widehat{w}_{\mathbf{G}}\in\operatornamewithlimits{argmin}_{w\in\mathbb{R}^{p}}H_{\ell}(w,\mathbf{G}):=\operatornamewithlimits{argmin}_{w\in\mathbb{R}^{p}}\left\{\frac{1}{n}\sum_{i=1}^{n}\ell_{i}(G_{i}^{\top}w)+R(w)\right\}.$ (6)
Note that numerous studies has investigated the asymptotic behaviour of the estimator and the risk in the proportionally high-dimensional limit.
3.2. Universality of Empirical Risk
We show the universality of the minimized empirical risk under $d$-block dependence. Specifically, we show that the distributions of $\min_{w}H_{\ell}(w,\mathbf{X})$ and $\min_{w}H_{\ell}(w,\mathbf{G})$ are asymptotically equivalent. The following theorem evaluates the distance between the two distributions under general $d$, using any $C^{3}(\mathbb{R})$ class function $g$:
Theorem 1.
Suppose Assumptions 1, 2 and 3 hold. Also, suppose that the distribution of $X_{1},...,X_{n}$ is $d$-block dependent. Then, for sufficiently large $n$, there exists some $C_{0}=C_{0}(\tau,q,M)>0$ such that the following hold: for any $\mathcal{S}_{n}\subset[-L_{n},L_{n}]^{p}$ with $L_{n}\geq 1$, and any $g\in C^{3}(\mathbb{R})$, we have
$\left|E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{X})\right)\right]-E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{G})\right)\right]\right|\leq C_{0}\cdot K_{g}\cdot r_{R}(L_{n}).$
Here $K_{g}:=1+\max_{\ell\in[0:3]}\|g^{(\ell)}\|_{\infty}$ and $r_{R}(L_{n})$ is defined by
$r_{R}(L_{n}):=\inf_{\rho\in(0,\bar{\rho})}\left\{\mathcal{M}_{\ell}(\rho)+\mathcal{D}^{3}_{\ell}(\rho)\inf_{\delta\in(0,\omega_{n})}\left[\mathcal{M}_{R}(L_{n},\delta)+Av(\{L_{\ell_{i}}^{\mathbf{q}}\})^{1/3}\cdot L_{n}^{(\bar{\mathbf{q}}_{0}+3)/{3}}\log^{2/3}_{+}(L_{n}/\delta)\sigma^{1/3}_{n}\right]\right\}$
where we define $\omega_{n}:=n^{-(q_{0}q_{1}+3)/2}\sigma_{n}$ and
$\sigma_{n}:=\left(\frac{pd^{3+(\bar{\mathbf{q}}_{0}/2)}}{n^{5/2}}+\frac{pd^{2}}{n^{3/2}}+\frac{pd^{\bar{\mathbf{q}}_{0}+2}}{n^{(\bar{\mathbf{q}}_{0}+3)/2}}\right).$ (7)
Consequently, for any $z\in\mathbb{R}$, $\varepsilon>0$, we obtain
$\mathbb{P}\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{X})>z+3\varepsilon\right)\leq\mathbb{P}\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{G})>z+\varepsilon\right)+C_{1}(\max\{1,\varepsilon^{-3}\})r_{R}(L_{n}).$
and
$\mathbb{P}\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{G})>z+3\varepsilon\right)\leq\mathbb{P}\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{X})>z+\varepsilon\right)+C_{1}(\max\{1,\varepsilon^{-3}\})r_{R}(L_{n}).$
Here $C_{1}>0$ is an absolute multiple of $C_{0}$.
Theorem 1 shows that the difference in expectations of the minimized risk functions is characterized by $r_{R}(L_{n})$, which includes the properties of the loss function $\ell_{i}$ and the regularizer $R$. The asymptotic equivalence of the distributions is guaranteed if $r_{R}(L_{n})$ converges to to zero, which largely on the convergence of $\sigma_{n}$.
From the result of Theorem 1, we can discuss the value of $d$ required to converge the error to zero. The following corollary summarizes the result:
Corollary 2.
Consider one of the loss functions presented in Example 1. Suppose that the conditions in Theorem 1, $\bar{\mathbf{q}}_{0}\leq 6$, and $Av(\{L_{\ell_{i}}^{\mathbf{q}}\})=O(1)$ hold. Set $L_{n}=O(n^{\alpha}\log^{\kappa}n)$ with $\alpha\in[0,1/(2\bar{\mathbf{q}}_{0}+6)]$ and $\kappa\geq 0$ as $n\to\infty$. Then, with setting $d=o(n^{1/4-(\bar{\mathbf{q}}_{0}+3)\alpha/2}/\log^{6+(2\bar{\mathbf{q}}_{0}+6)\kappa}n)$ as $n\to\infty$, for any $\mathcal{S}_{n}\subset[-L_{n},L_{n}]^{p}$ with $L_{n}\geq 1$, and any $g\in C^{3}(\mathbb{R})$, we have the following as $n,p\to\infty$:
$\left|E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{X})\right)\right]-E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{G})\right)\right]\right|\to 0.$
We first note the setup of $L_{n}$. As seen from Section 3 in Han and Shen (2023) or Section 7 in Lahiry and Sur (2024), typical estimator examples satisfy $\bar{\mathbf{q}}_{0}\leq 6$, and also $L_{n}=O(\log^{\kappa}n)$ with $\kappa\geq 0$ is enough to cover a support of distributions of the estimators. In this situation, we can set $\alpha=0$ and obtain an order $d=o(n^{1/4})$ up to logarithmic factors that makes the discrepancy converge to zero in the proportionally high-dimensional limit.
We discuss the relationship between Theorem 1 and previous results of the existing studies. Theorem 1 can be regarded as a generalization of Theorem 2.3 in Han and Shen (2023). Han and Shen (2023) considers the case where all elements of $X_{i}$ are independent, i.e., $d$-block dependent with $d=1$, and the statement of their theorem is equivalent to Theorem 1 with $\sigma_{n}$ replaced by $n^{-1/2}$. Additionally, Theorem 4.1 of Lahiry and Sur (2024) has a similar universality statement under the block dependence, where $d=o((n/\log n)^{1/5})$ holds. Note that the permissible order of Lahiry and Sur (2024) is more restrictive than those allowed in Theorem 1 and Corollary 2 with $\alpha=0$. Furthermore, Chang et al. (2024) shows the universality of a maximum of high-dimensional random vector with the dependent random variables. Although there is some technical similarity, the result by Chang et al. (2024) is not comparable with our result, because their setup is different from our setting. For details, see Remark 2.
3.3. Universality of Estimator
We show the universality of the estimator, which is the optimizer of the empirical risk. Specifically, we demonstrate that the estimator $\widehat{w}_{\mathbf{X}}$ defined in (3) and the estimator $\widehat{w}_{\mathbf{G}}$ under the Gaussian covariate asymptotically have the same distribution.
Theorem 3.
Suppose Assumptions 1, 2 and 3 hold. Also, suppose that the distribution of $X_{1},...,X_{n}$ is $d$-block dependent. Fix a measurable subset $\mathcal{S}_{n}\subset\mathbb{R}^{p}$. Suppose there exist $z\in\mathbb{R},\rho_{0}>0$, $L_{n}\geq 1$ and $\varepsilon_{n}\in[0,1/4)$ such that the following hold:
  1. (i)
    Both $\|\widehat{w}_{\mathbf{G}}\|_{\infty}$ and $\|\widehat{w}_{\mathbf{X}}\|_{\infty}$ increase moderately in the sense that
    $\max\left\{\mathbb{P}(\|\widehat{w}_{\mathbf{G}}\|_{\infty}>L_{n}),\mathbb{P}(\|\widehat{w}_{\mathbf{X}}\|_{\infty}>L_{n})\right\}\leq\varepsilon_{n}$
  2. (ii)
    $\widehat{w}_{\mathbf{G}}$ does not satisfy the specific structure $\mathcal{S}_{n}$ in the sense that
    $\max\left\{\mathbb{P}\left(\min_{w\in\mathbb{R}^{p}}H_{\ell}(w,\mathbf{G})\geq z+\rho_{0}\right),\mathbb{P}\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{G})\leq z+2\rho_{0}\right)\right\}\leq\varepsilon_{n}.$
Then, $\widehat{w}_{\mathbf{X}}$ also does not satisfy the specific structure $\mathcal{S}_{n}$ with high probability:
$\mathbb{P}\left(\widehat{w}_{\mathbf{X}}\in\mathcal{S}_{n}\right)\leq 4\varepsilon_{n}+C_{0}\max\{1,\rho_{0}^{-3}\}r_{R}(L_{n})$
Here, $r_{R}(L_{n})$ is defined in Theorem 1, and $C_{0}>0$ depends on $\tau,q,M$ only.
This theorem shows that if $r_{R}(L_{n})$ converges to zero under the conditions given in Theorem 1, the distributions of the two estimators become asymptotically equivalent. As discussed in Section 3.2, the convergence of $r_{R}(L_{n})$ is ensured when $d=o(n^{1/4})$ holds up to logarithmic factors. It is summarized in the following statement:
Corollary 4.
Consider one of the loss functions presented in Example 1. Suppose that the conditions of Theorem 3 hold with a fixed $\mathcal{S}_{n}\subset\mathbb{R}^{p}$ and $\varepsilon_{n}\in[0,1/4)$. $\bar{\mathbf{q}}_{0}\leq 6$, and $Av(\{L_{\ell_{i}}^{\mathbf{q}}\})=O(1)$ hold. Set $L_{n}=O(n^{\alpha}\log^{\kappa}n)$ with $\alpha\in[0,1/(2\bar{\mathbf{q}}_{0}+6)]$ and $\kappa\geq 0$ as $n\to\infty$. Then, with setting $d=o(n^{1/4-(\bar{\mathbf{q}}_{0}+3)\alpha/2}/\log^{6+(2\bar{\mathbf{q}}_{0}+6)\kappa}n)$ as $n\to\infty$, we have the following as $n,p\to\infty$:
$|\mathbb{P}\left(\widehat{w}_{\mathbf{X}}\in\mathcal{S}_{n}\right)-\mathbb{P}\left(\widehat{w}_{\mathbf{G}}\in\mathcal{S}_{n}\right)|\leq 4\varepsilon_{n}+o(1).$
We provide explanations on conditions (i) and (ii) in this Theorem 3. Condition (i) requires the boundedness of the estimators $\widehat{w}_{\mathbf{X}}$ and $\widehat{w}_{\mathbf{G}}$, which is demonstrated depending on each example. Condition (ii) can be investigated by a limiting value of the minimum of the empirical risk in the proportionally high-dimension regime with Gaussian covariate, which also depends on specific examples. These two conditions are identical to those provided in Theorem 2.4 of Han and Shen (2023), and more detailed discussions will be given in the subsequent section on applications.
4. Outline of Proof with Lindeberg Principle
In this section, we outline the proof of Theorem 1. To emphasize the fact that Theorem 1 still holds even though the Gaussianity of $\mathbf{G}$ is violated, we replace $\mathbf{G}$ with another design matrix $\mathbf{W}=(W_{1},...,W_{n})^{\top}\in\mathbb{R}^{n\times p}$ that has the same first and second moments as $\mathbf{X}$, that is, $\mathbb{E}[W_{i}]=\mathbb{E}[X_{i}]=0$ and $\mathbb{E}[W_{i}W_{i}^{\top}]=\mathbb{E}[X_{i}X_{i}^{\top}]=\Sigma/n$ hold for $i=1,...,n$.
We prepare an additional notation to present the outline. We define a smoothly approximated version of the empirical risk as $H_{\ell_{\rho}}(w,\mathbf{X}):=\frac{1}{n}\sum_{i=1}^{n}\ell_{i,\rho}(X_{i}^{\top}w)+R(w)$. Let $\bar{H}_{\ell}$ denote the unnormalized function $H_{\ell}$, namely $n\cdot H_{\ell}(w,\mathbf{X})$. In addition, we define a finite subset $\mathcal{S}_{n,\delta}\subset\mathcal{S}_{n}$ depending on $\delta>0$, which will be specified later. Furthermore, we introduce the soft minimum function that approximates the minimum function: for $\beta>0$ and $N\in\mathbb{N}$, we define the soft minimum function $F_{\beta}$ as
$F_{\beta}(x):=-\frac{1}{\beta}\log\left(\sum_{j=1}^{N}\exp(-\beta x_{j})\right),\quad x\in\mathbb{R}^{N}.$
4.1. Overview
We focus on the difference $E[g(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{X}))]-E[g(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{W}))]$ with $g\in C^{3}(\mathbb{R})$, which is bounded in Theorem 1 with replacing $\mathbf{G}$ by $\mathbf{W}$. The difference in expectations can be divided into four terms:
$\left|E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{X})\right)\right]-E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{W})\right)\right]\right|$
$\leq 2\max_{\mathbf{Z}\in\{\mathbf{X},\mathbf{W}\}}\left|E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{Z})\right)\right]-E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell_{\rho}}(w,\mathbf{Z})\right)\right]\right|$ (8)
$\quad+2\max_{\mathbf{Z}\in\{\mathbf{X},\mathbf{W}\}}\left|E\left[g\left(\min_{w\in\mathcal{S}_{n,\delta}}H_{\ell_{\rho}}(w,\mathbf{Z})\right)\right]-E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell_{\rho}}(w,\mathbf{Z})\right)\right]\right|$ (9)
$\quad+2\max_{\mathbf{Z}\in\{\mathbf{X},\mathbf{W}\}}\left|E\left[g\left(\min_{w\in\mathcal{S}_{n,\delta}}H_{\ell_{\rho}}(w,\mathbf{Z})\right)\right]-E\left[g\left(n^{-1}F_{\beta}\left((\bar{H}_{\ell_{\rho}}(w,\mathbf{Z}))_{w\in\mathcal{S}_{n,\delta}}\right)\right)\right]\right|$ (10)
$\quad+\left|E\left[g\left(n^{-1}F_{\beta}\left((\bar{H}_{\ell_{\rho}}(w,\mathbf{X}))_{w\in\mathcal{S}_{n,\delta}}\right)\right)\right]-E\left[g\left(n^{-1}F_{\beta}\left((\bar{H}_{\ell_{\rho}}(w,\mathbf{W}))_{w\in\mathcal{S}_{n,\delta}}\right)\right)\right]\right|$ (11)
$=:I+II+III+IV.$ (12)
where we define $\bar{H}=n\times H$. The terms $I,II,III$, and $IV$ each have their own intuition. $I$ is on a smooth approximation error of the loss function by $\ell_{\rho}$, $II$ is about the discretization on the parameter space from $\mathcal{S}_{n}$ to $\mathcal{S}_{n,\delta}$, $III$ is an approximation error by the soft minimum function $F_{\beta}(\cdot)$, and $IV$ is an effect of approximating the function via moment matching technique.
Our technical novelty lies in studying $IV$ with the block dependency. Specifically, we develop the Lindeberg-type approach to derive an upper bound deriving an upper bound for $IV$ in terms of $d$. The remaining terms $I,II$, and $III$ are evaluated by the proof developed in Han and Shen (2023), respectively. We first state the proof outline for studying the term $IV$ and then explain how to bound the remaining terms $I,II$, and $III$.
4.2. Bound for $IV$: Lindeberg Principle with Block Dependence
In this section, we derive an approximation result for $d$-block dependent random variables $X_{i}$ and $W_{i}$. This proof employs two crucial lemmas. The first lemma is a generalization of the Lindeberg principle, applicable to $d$-block dependent random variables. The second lemma provides an upper bound for the sum of $\sum_{j=1}^{k}|B_{j}|^{m}$ for $m\geq 0$. Given that the sum of $|B_{i}|$ equals the dimension of covariates, we show that, for any $m\in\mathbb{N}$, the sum of $|B_{i}|^{m}$ can be bounded above by a term of order $pd^{m-1}$ with two key techniques (i) block replacement, and (ii) construction of the aligned blocks.
4.2.1. Generalized Lindeberg Principle
We develop a generalized Lindberg principle, which is a method used in evaluating discrepancy between random vectors by swapping their elements and evaluating the error using the Taylor approximation. Specifically, we employ this principle to evaluate the discrepancy of random matrices with block dependence. The independence of $B_{j}$ from any other $B_{i}$ allows us to upper bound the difference in expectations for any function $f\in C^{3}(\mathbb{R}^{n\times p})$.
We provide a key result of the Lindeberg principle with the block dependence. We define $p$-dimensional random vector $X_{i}^{(B_{k})}:=(X_{i,1}\textbf{1}\{1\in B_{k}\},...,X_{i,p}\textbf{1}\{p\in B_{k}\})^{\top}$. Then, we obtain the following lemmas.
Lemma 5.
Consider two sequences of i.i.d. $\mathbb{R}^{p}$-valued random vectors $X_{1},...,X_{n}$ and $W_{1},...,W_{n}$. Suppose that each $\{X_{i}\}_{i=1}^{n}$ and $\{W_{i}\}_{i=1}^{n}$ is $d$-block dependent, and $E[X_{i}]=E[W_{i}]$ and $E[X_{i}X_{i}^{\top}]=E[W_{i}W_{i}^{\top}]$ hold. Then, for any $f\in C^{3}(\mathbb{R}^{n\times p})$ satisfying $\max_{i\in[n],j\in[p]}|E[\partial_{i,j}f(\mathbf{Z}^{0}_{i,j})]|<\infty$ and $\max_{i\in[n],j,k\in[p]}|E[\partial_{i,j}\partial_{i,k}f(\mathbf{Z}^{0}_{i,j})]|<\infty$, we have
$|E[f(\mathbf{X})]-E[f(\mathbf{W})]|$
$\leq\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{k}\left|E\left[\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})dt\right]\right|$
$~{}~{}+\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{k}\left|E\left[\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}W_{i,\ell}W_{i,\ell^{\prime}}W_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tW^{(B_{j})}_{i})dt\right]\right|,$
where
$\mathbf{Z}_{i,j}^{0}$ $:=\left(X_{1},\cdots,X_{i-1},(W_{i}-X_{i}^{(B_{j})}+\sum_{\ell=1}^{j}(X_{i}^{(B_{\ell})}-W_{i}^{(B_{\ell})})),W_{i+1},\cdots,W_{n}\right)^{\top}.$
This result indicates that the difference in the distributions of $\mathbf{X}$ and $\mathbf{W}$ is evaluated based on the size of $B_{j}$ of the block dependencies. Since there is a triple summation over $B_{j}$ within the upper bound of the lemma, the cubic value $|B_{j}|^{3}$ becomes critical. If we have $|B_{j}|=1$, which corresponds to the case where all elements of $X_{i}$ and $W_{i}$ are independent, the expectation terms of the upper bound remains bounded. In the following sections, it becomes crucial to assess $\sum_{j}|B_{j}|^{m}$ with $m\geq 0$ so that this upper bound does not diverge.
4.2.2. Bounding Sum of Block Size
In this section, we provide an upper bound for the sum of the powers of the size of cells $\sum_{j=1}^{k}|B_{j}|^{m}$. This bound is useful when evaluating the sum of the cubic values $|B_{j}|^{3}$ introduced in the previous section. The result and the proof are presented below:
Lemma 6.
Consider the $k$-partition $\{B_{j}\}_{j=1}^{k}$ of $[p]$, which is a sequence of $k$ cells $B_{1},...,B_{k}\subset[p]$ that are disjoint to each other and satisfy $\cup_{j=1}^{k}B_{j}=[p]$. Also, suppose that $\max_{j=1,...,k}|B_{j}|\leq d$ holds with $1\leq d\leq p$. Then, for any $m\in\mathbb{N}$, we have
$\sum_{j=1}^{k}|B_{j}|^{m}\leq 4pd^{m-1}.$ (13)
Proof.
In the case of $m=1$, $\sum_{j=1}^{k}|B_{j}|=p$ holds by its definition. Hence, we obtain the statement.
In the case of $m\geq 2$, to obtain the statement, we generate newly aligned cells $D_{1},...,D_{s}\subset[p]$ from the original cells of the partition $\{B_{j}\}_{j=1}^{k}$ by the following steps described by a pseudo code:
  1. 1.
    Define $I:=\{\ell\in[k]\mid|B_{\ell}|\geq\lfloor d/2\rfloor+1\}$ and $I^{c}:=[k]\backslash I$.
  2. 2.
    Initialize $i\leftarrow 1$.
  3. 3.
    For $j\in I$, repeat the followings:
    1. i.
      Define $D_{i}:=B_{j}$.
    2. ii.
      Set $i\leftarrow i+1$.
  4. 4.
    For $j\in I^{c}$, repeat the followings:
    1. i.
      Define $D_{i}:=B_{j}$.
    2. ii.
      Set $I^{c}\leftarrow I^{c}\backslash\{j\}$.
    3. iii.
      If $I^{c}=\emptyset$, then break.
    4. iv.
      For $j^{\prime}\in I^{c}$, repeat the followings:
      1. a.
        Set $I^{c}\leftarrow I^{c}\backslash\{j^{\prime}\}$.
      2. b.
        Set $D_{i}\leftarrow D_{i}\cup B_{j^{\prime}}$.
      3. c.
        If $|D_{i}|\geq\lfloor d/2\rfloor+1$ or $I^{c}=\emptyset$, then break.
    5. v.
      If $I^{c}=\emptyset$, then break:
    6. vi.
      Set $i\leftarrow i+1$.
  5. 5.
    Define $s\leftarrow i$.
  6. 6.
    Output $D_{1},D_{2},...,D_{s}$.
From the steps, the generated cells $D_{1},D_{2},\cdots,D_{s}\subset[p]$ satisfy $\sum_{j=1}^{k}|B_{j}|^{m}\leq\sum_{\ell=1}^{s}|D_{\ell}|^{m}$. Furthermore, it is clear from the algorithm that at most one cell $D_{j}$ satisfies $|D_{j}|\leq\lfloor d/2\rfloor$, while all the other cells satisfy $|D_{j}|\geq\lfloor d/2\rfloor+1$. Hence, $s$ satisfies the following inequality:
$s$ $\leq 1+\lceil p/(\lfloor d/2\rfloor+1)\rceil\leq 1+\lceil 2p/d\rceil\leq 2p/d+2\leq 4(p/d),$
where the last inequality holds from the assumptions that $d\leq p$. Therefore, we have
$\sum_{j=1}^{k}|B_{j}|^{m}$ $\leq\sum_{\ell=1}^{s}|D_{\ell}|^{m}\leq\sum_{\ell=1}^{s}d^{m}\leq 4pd^{m-1}.$
The important consequence in Lemma 6 is that the upper bound holds independently of $k$. In other words, as long as we determine the upper bound $d$ of the block size, we do not need to care how many blocks are used to divide the $p$ variables. In the proof, we arrange the existing cells to construct new cells, which are then used to control the total sum.
4.2.3. Unifying These Results
Using Lemma 5 and Lemma 6 above, we bound the term $IV$, which describes the difference between the distributions of the minimized risk with the datasets. The statement follows $g\in C^{3}(\mathbb{R}^{n\times p})$.
Proposition 7.
Suppose Assumptions 1, 2, and 3 hold. Consider two sequences of i.i.d. $\mathbb{R}^{p}$-valued random vectors $X_{1},...,X_{n}$ and $W_{1},...,W_{n}$. Suppose that each $\{X_{i}\}_{i=1}^{n}$ and $\{W_{i}\}_{i=1}^{n}$ is $d$-block dependent, and $E[X_{i}]=E[W_{i}]$ and $E[X_{i}X_{i}^{\top}]=E[W_{i}W_{i}^{\top}]$ hold. Then, there exists some $C_{0}=C_{0}(\tau,q,M)>0$ such that for any finite set $\mathcal{W}\subset[-L_{n},L_{n}]^{p}$ with $L_{n}\geq 1$, any $\rho\in(0,1)$, and any $g\in C^{3}(\mathbb{R}^{n\times p})$ satisfying $\max_{i\in[n],j\in[p]}|E[\partial_{i,j}g(\mathbf{Z}^{0}_{i})]|<\infty$ and $\max_{i\in[n],j,k\in[p]}|E[\partial_{i,j}\partial_{i,k}g(\mathbf{Z}^{0}_{i})]|<\infty$, we have
$|E[g(n^{-1}F_{\beta}((\bar{H}_{\ell_{\rho}}(w,\mathbf{X})))_{w\in\mathcal{W}}]-E[g(n^{-1}F_{\beta}((\bar{H}_{\ell_{\rho}}(w,\mathbf{W})))_{w\in\mathcal{W}}]|$
$\leq C_{0}(\tau,q,M)K_{g}\max\{1,\beta^{2}\}L_{n}^{\bar{\mathbf{q}}_{0}+3}D^{3}_{\ell}(\rho)Av(\{L^{\mathbf{q}}_{\ell_{i}}\})\sigma_{n}.$
By setting the arbitrarily set $\mathcal{W}$ here as $\mathcal{S}_{n,\delta}$, we develop an upper bound for the term $IV$. Note that the definition of $\sigma_{n}$ is given in the statement of Theorem 1.
4.3. Bound for Other Terms
We examine the other remaining items. Although the results here are the same as in Han and Shen (2023), we explain their outline for the sake of completeness.
4.3.1. Bound for $I$: Smooth Approximation
We bound the term $I$ on the smooth approximation for a loss function $\ell_{i}$. From Assumption 2, $\|\ell_{i}-\ell_{i;\rho}\|_{\infty}$ is upper bounded by $\mathcal{M}_{\ell}(\rho)$. Hence, for any $\mathbf{Z}\in\{\mathbf{X},\mathbf{W}\}$, we naturally obtain
$\left|E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{Z})\right)\right]-E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell_{\rho}}(w,\mathbf{Z})\right)\right]\right|$
$\leq\|g^{\prime}\|_{\infty}E\left[\left|\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{Z})-\min_{w\in\mathcal{S}_{n}}H_{\ell_{\rho}}(w,\mathbf{Z})\right|\right]$
$\leq K_{g}\mathcal{M}_{\ell}(\rho).$
4.3.2. Bound for $II$: Discretization of Parameter Space
We study the term $II$, which explains an effect of approximating the minimization problem $\min_{w\in\mathcal{S}_{n}}H_{\ell{\rho}}(w,\mathbf{X})$ on the parameter space $\mathcal{S}_{n}$ to the minimization problem $\min_{w\in\mathcal{S}_{n,\delta}}H_{\ell{\rho}}(w,\mathbf{X})$ on a discretized parameter space $\mathcal{S}_{n,\delta}$. To the aim, we obtain the following statement.
Proposition 8.
Suppose Assumption 2 holds. Fix $L_{n}\geq 1$. There exists some $C_{0}=C_{0}(q)>0$, such that for any measurable $\mathcal{S}_{n}\subset[-L_{n},L_{n}]^{p}$ and any $\delta\in(0,1)$, there exists a deterministic finite set $\mathcal{S}_{n,\delta}$ with $|\mathcal{S}_{n,\delta}|\leq\lceil 2L_{n}/\delta\rceil^{p}$ and the following hold: for any $\rho\in(0,\bar{\rho})$, $\mathbf{Z}\in\{\mathbf{X},\mathbf{W}\}$ and $g\in C^{1}(\mathbb{R})$, we have
$\left|E\left[g\left(\min_{w\in\mathcal{S}_{n,\delta}}H_{\ell_{\rho}}(w,\mathbf{Z})\right)\right]-E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell_{\rho}}(w,\mathbf{Z})\right)\right]\right|$
$\leq C_{0}\|g^{\prime}\|_{\infty}\left(\mathcal{D}_{\ell}(\rho)\delta\cdot L_{n}^{q_{0}q_{1}}Av({L_{\ell_{i}}^{q_{1}}})\cdot\left(E\left[\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|\right]+E\left[\left(\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|\right)^{q_{0}q_{1}+1}\right]\right)+\mathcal{M}_{R}(L_{n},\delta)\right).$
This result controls the error that arises when optimization is performed over a discretized finite set $\mathcal{S}_{n,\delta}$ instead of the original parameter set $\mathcal{S}_{n}$. Although this result is similar to Equation (4.11) in Han and Shen (2023), there are several differences. By taking $\delta$ sufficiently small, the upper bound of this error can be made arbitrarily small. The smoothness of a loss function and a regularizer allows us to achieve this result. By assigning each value in $\mathcal{S}_{n;\delta}$ to a loss function, we can represent them as a high-dimensional vector. Let $(H(w,\mathbf{X}))_{w\in\mathcal{S}_{n;\delta}}$ denote this vector with covariate $\mathbf{X}$. This step plays an important role as reducing the approximation problem to approximating two high-dimensional vectors: $(H(w,\mathbf{X}))_{w\in\mathcal{S}_{n;\delta}}$ and $(H(w,\mathbf{W}))_{w\in\mathcal{S}_{n;\delta}}$.
4.3.3. Bound for $III$: Smooth minimum approximation
We study the term $III$ to evaluate the approximation error by the smooth minimum function $F_{\beta}$. A straightforward calculation yields approximation to minimum function as
$\min_{1\leq j\leq N}x_{j}-\frac{\log N}{\beta}\leq F_{\beta}(x)\leq\min_{1\leq j\leq N}x_{j},~{}x\in\mathbb{R}^{N}.$
By utilizing this property, we obtain the following proposition.
Proposition 9.
For any finite set $\mathcal{W}\subset\mathbb{R}^{p}$ and $g\in C^{1}(\mathbb{R})$, we have
$\left|E\left[g\left(\min_{w\in\mathcal{W}}H_{\ell_{\rho}}(w,\mathbf{X})\right)\right]-E\left[g\left[n^{-1}F_{\beta}\left((\bar{H}_{\ell_{\rho}}(w,\mathbf{X}))_{w\in\mathcal{W}}\right)\right]\right]\right|\leq\frac{\|g^{{}^{\prime}}\|_{\infty}\log|W|}{n\beta}$
Evaluation by the soft minimum function is one of the useful tools used in high-dimensional analysis, as demonstrated by Chernozhukov et al. (2013). By exploiting the smoothness of the soft minimum function, we can apply various approximation methods, such as Stein’s method and the generalized Lindeberg principle.
Remark 2 (Connection to high-dimensional Gaussian approximation studies).
In this proof, we employ several techniques used in the high-dimensional Gaussian approximation studies (Chernozhukov et al., 2013, 2014; Chernozhuokov et al., 2022; Lopes, 2022; Chang et al., 2024). However, not all techniques from the approximation can be applied directly. There are two main reasons for this. First, there is a difference in the objects of differentiation when considering neighborhood approximations. In the approximation studies, one needs to differentiate the smoothly approximated maximum of a high-dimensional data vector with respect to the high-dimensional data itself. In contrast, in our setting, the loss function maximized with respect to the parameters is differentiated by the values of the high-dimensional data. Since the values being maximized and differentiated differ in this way, we cannot apply the approximation studies in our setup. Second, the use of Nazarov’s inequality plays a crucial role. While some of the approximation study Chang et al. (2024) needs Nazarov’s inequality, Nazarov’s inequality also does not hold for the design matrix $\mathbf{X}$ or $\mathbf{W}$, since we consider the explicit form of the minimum of the empirical risk function. Due to these differences, the results from the studies on high-dimensional Gaussian approximations cannot be directly applied.
5. Application to Distribution of Robust Estimators
In this section, we demonstrate the universality of the distribution of estimators with block dependence under specific loss functions designed for robustness. Specifically, we consider estimators under the absolute loss and the Huber loss, as presented in Example 1. We define a robust cost function as
$H^{R}(w,\mathbf{X},\xi):=\frac{1}{n}\sum_{i=1}^{n}\ell_{0}((X_{i}^{\top}w)_{i}-\xi_{i})+\frac{\lambda}{2}\|w+\theta_{0}\|^{2}.$
where $\lambda>0$ is a coefficient for the regularizer. Let $\widehat{w}^{R}_{\mathbf{X}}$ denote the minimizer of the cost function with covariates $\mathbf{X}$. Recall that $\theta_{0}\in\mathbb{R}^{p}$ be the true parameter of the regression model (1), and the estimator of the regression model as the minimizer (2) with the setup is given as $\widehat{\theta}_{\mathbf{X}}^{R}=\widehat{w}_{\mathbf{X}}^{R}+\theta_{0}$. By substituting a Gaussian design matrix $\mathbf{G}$ instead of $\mathbf{X}$, we can consider the Gaussian analogue $H^{R}(w,\mathbf{G},\xi)$ as Section 3.1.
To show the universality, we introduce the following assumptions:
Assumption 4 (Assumptions on Robust estimators).
The followings hold:
  1. i.
    $\ell_{0}:\mathbb{R}\rightarrow\mathbb{R}$ is convex with a weak derivative $\ell^{\prime}_{0}$ satisfying $\max\{\ell_{0}(0),\operatorname*{ess\,sup}|\ell^{\prime}_{0}|\}\leq L_{0}$ for some $L_{0}>0$.
  2. ii.
    There exists $0<\varepsilon\leq(1/4)$ such that $d=o(n^{{1}/{4}-\varepsilon})$
  3. iii.
    A rescaled covariate $X^{0}_{i}=\sqrt{n}X_{i}$ satisfies $M_{k^{*}:\mathbf{X}}:=\max_{j\in[p]}E[|X^{0}_{i,j}|^{2\lceil k^{*}/2\rceil}]<\infty$ where $k^{*}>{3}/({2\varepsilon})$. Also, $\lambda_{\max}(\Sigma)<\infty$ holds.
Condition i generalizes the condition previously mentioned for the robustness of the loss. Both the absolute loss and the Huber loss satisfy this condition. By combining Conditions ii and iii, we can explain how the moment condition of the scaled covariates $X_{i}^{0}$ is affected by the order of $d$ for the block dependence.
To apply the result of Theorem 3, we establish the upper bound for the moment of robust estimators.
Lemma 10.
Suppose Assumptions 1 and 4 hold. Also, suppose that the distribution of $X_{1},...,X_{n}$ is $d$-block dependent. Then, for any $k\geq 2$, we have $K:=K(p)>0$ satisfying
$\max_{j\in[p]}E\left[\left|\widehat{w}^{R}_{\mathbf{X},j}\right|\right]\leq K\cdot\left\{(L_{0}/\lambda)^{k}(M_{k:\mathbf{X}})^{k/(2\lceil k/2\rceil)}+\|\theta_{0}\|^{k}_{\infty}\right\}.$
By combining the result with Theorem 3, we obtain the following universality result on the distribution of the estimator $\widehat{w}^{R}_{\mathbf{X}}$:
Theorem 11.
Suppose Assumptions 1 and 4 hold. Also, suppose that the distribution of $X_{1},...,X_{n}$ is $d$-block dependent. Fix a measurable subset $\mathcal{S}_{n}\subset\mathbb{R}^{p}$. Suppose that there exist $z\in\mathbb{R},\rho_{0}>0$, and $\varepsilon_{n}\in[0,1/4)$ such that the following hold:
  1. (ii)
    $\widehat{w}_{\mathbf{G}}$ does not satisfy the specific structure $\mathcal{S}_{n}$ in the sense that
    $\max\left\{\mathbb{P}\left(\min_{w\in\mathbb{R}^{p}}H^{R}(w,\mathbf{G},\xi)\geq z+\rho_{0}\right),\mathbb{P}\left(\min_{w\in\mathcal{S}_{n}}H^{R}(w,\mathbf{G},\xi)\leq z+2\rho_{0}\right)\right\}\leq\varepsilon_{n}.$
Then, $\widehat{w}_{\mathbf{X}}$ also does not satisfy the specific structure $\mathcal{S}_{n}$ with high probability, i.e. there exists some $K=K(\lambda,\tau,M_{k^{*}:\mathbf{X}},k^{*},\varepsilon,L_{0})>0$ such that
$\mathbb{P}(\widehat{w}^{R}_{\mathbf{X}}\in\mathcal{S}_{n})\leq 4\varepsilon_{n}+K(1+\|\theta_{0}\|^{k^{*}}_{\infty}+\rho^{-3}_{0})n^{-{1}/{45}}$
With the result the universality of the distribution, Theorem 11 will derive the convergence result of estimation error with non-Gaussian covariates:
Theorem 12.
Suppose that Assumption 1 and 4 hold with $\sup_{p}\max_{j\in[p]}E[|X^{0}_{i,j}|^{2\lceil k^{*}/2\rceil}]<\infty$, and define $\tau_{0}=(n/p)$. Moreover. we assume the followings:
  1. (i)
    Either (a) $\ell_{0}$ does not have derivative at certain point, or (b) $\ell_{0}$ includes an interval where $\psi_{0}$ has a strictly increasing derivative.
  2. (ii)
    A distribution of $\xi$ is absolutely continuous with respect to the Lebesgue measure.
  3. (iii)
    $\theta_{0}$ contains i.i.d. components whose law $\Pi_{0}$ guarantees the existence of moment of any order.
Then with $Z\sim N(0,1)$, the system of equations
$(\gamma^{R}_{*})^{2}/\tau_{0}=$ $E\left[\gamma^{R}_{*}Z+\xi_{1}-\mathsf{prox}_{\ell_{0}}\left(\gamma^{R}_{*}Z+\xi_{i};\beta^{R}_{*}\right)\right]+\lambda^{2}(\beta^{R}_{*})^{2}\cdot E\Pi^{2}_{0},$
$1-\tau_{0}^{-1}+\lambda\beta^{R}_{*}=$ $E\left[\mathbf{\mathsf{prox}^{\prime}}_{\ell_{0}}\left(\gamma^{R}_{*}Z+\xi_{i};\beta^{R}_{*}\right)\right]$
have a unique solution $(\beta^{R}_{*},\gamma^{R}_{*})\in(0,\infty)$ such that we have
[…]
This result demonstrates that the error of the estimator $\widehat{\theta}_{\mathbf{X}}$ form non-Gaussian covariates converges to the known value $\tau_{0}(\gamma^{R}_{*})^{2}$, which is given by the solution of the system. The specification of the system has been well studied in the context of Gaussian covariates, e.g. Thrampoulidis et al. (2015, 2018). Our result shows that this universality of the estimation error also holds when the independence assumption of covariates $X_{i}$ is relaxed.
Remark 3 (Possibility of extension to lasso/ridge regression).
We discuss extending this result to the Lasso and Ridge regressions under the squared loss, as examined in Han and Shen (2023). A key point here is whether the upper bound for the estimator holds, as in Lemma 10 in the setup in this section. In Han and Shen (2023), the upper bound is derived under the assumption that each element of $X_{i}$ is independent, so the bound does not apply to our case directly. Instead, we can utilize the upper bound under dependent covariates, as in Lemma 7.3 of Lahiry and Sur (2024), which allows us to show the universality of the Lasso and Ridge regressions.
6. Conclusion
We derive the universality of the estimator for linear models under the assumption that the distribution of covariates exhibits the block dependence. While the properties of estimators in many high-dimensional models are typically identified under the assumption of Gaussian covariates, extending these results to settings with non-Gaussian covariates remains an important challenge. While the existing studies have required independence among each element of the covariates, this study relaxes that assumption and demonstrates that universality holds even under the block dependence of covariates of a certain order. This allows us to precisely identify the robust estimator for non-Gaussian covariates with dependence among elements.
Appendix A Proof of Main Results
A.1. Proof of Theorem 1
Proof.
From Proposition 7, 8, and 9, for any fixed $\delta\in(0,1)$, we obtain
$\left|E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{X})\right)\right]-E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{W})\right)\right]\right|$
$\lesssim_{\tau,q,M}K_{g}$ $\Biggl{[}\mathcal{M}_{\ell}(\rho)+\mathcal{D}^{3}_{\ell}(\rho)$
×$\times\left\{L_{n}^{q_{0}q_{1}}Av({L_{\ell_{i}}^{q_{1}}})\cdot\max_{Z_{ij}\in\{X_{ij},W_{ij}\}}\left(E\left[\max_{1\leq i\leq n}\sum_{j=1}^{p}|Z_{i,j}|\right]+E\left[\left(\max_{1\leq i\leq n}\sum_{j=1}^{p}|Z_{i,j}|\right)^{q_{0}q_{1}+1}\right]\right)\delta\right.$
$\quad\left.+\mathcal{M}_{R}(L_{n},\delta)+\beta^{-1}\log(L_{n}/\delta)+\max\{1,\beta^{2}\}L_{n}^{\bar{\mathbf{q}}_{0}+3}Av(\{L^{\mathbf{q}}_{\ell_{i}}\})\sigma_{n}\right\}\biggr{]}.$
By trivial calculation, for each $\mathbf{Z}\in\{\mathbf{X},\mathbf{W}\}$, we have
$E\left[\max_{1\leq i\leq n}\sum_{j=1}^{p}|Z_{i,j}|\right]+E\left[\left(\max_{1\leq i\leq n}\sum_{j=1}^{p}|Z_{i,j}|\right)^{q_{0}q_{1}+1}\right]\lesssim_{\tau,q,M}n^{(q_{0}q_{1}+3)/2}.$
Therefore, by setting $0<\delta\leq n^{-(q_{0}q_{1}+3)/2}\sigma_{n}$, the upper bound becomes
$\mathcal{M}_{\ell}(\rho)+\mathcal{D}^{3}_{\ell}(\rho)\left\{\mathcal{M}_{R}(L_{n},\delta)+\beta^{-1}\log(L_{n}/\delta)+\max\{1,\beta^{2}\}L_{n}^{\bar{\mathbf{q}}_{0}+3}Av(\{L^{\mathbf{q}}_{\ell_{i}}\})\sigma_{n}\right\}.$
By minimizing the above term in terms of $\beta$, we obtain the required result.
For the second claim, take any function $g\in C^{3}(\mathbb{R})$ such that
$g(x):=\begin{cases}0&\text{if $x\leq 1$,}\\ 2&\text{if $x>3$.}\end{cases}$
With $\varphi(\mathbf{X}):=\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{X})$, for any $z>0,\varepsilon>0$, we have
$\mathbb{P}(\varphi(\mathbf{X})-z>3\varepsilon)\leq$ $\mathbb{P}[g((\varphi(\mathbf{X})-z)/\varepsilon)\geq 2]$
$\leq$ $\frac{1}{2}E[g((\varphi(\mathbf{X})-z)/\varepsilon)]$
$\leq$ $\frac{1}{2}E[g((\varphi(\mathbf{W})-z)/\varepsilon)]+C\cdot\max\{1,\varepsilon^{-3}\}r_{R}(L_{n})$
$\leq$ $\mathbb{P}(\varphi(\mathbf{W})-z)>\varepsilon)+C\cdot\max\{1,\varepsilon^{-3}\}r_{R}(L_{n}).$
A.2. Proof of Theorem 3
Proof.
From the condition (1), for the event $A:=\{\widehat{w}_{\mathbf{X}}\in[-L_{n},L_{n}]^{p}\}$, we have $\mathbb{P}(A^{c})\leq\varepsilon_{n}$. Because the event
$\left\{\min_{\|w\|_{\infty}\leq L_{n}}H_{\ell}(w,\mathbf{X})<z+3\rho_{0}\right\}\bigcap\left\{\min_{w\in\mathcal{S}_{n}\cap[-L_{n},L_{n}]^{p}}H_{\ell}(w,\mathbf{X})>z+6\rho_{0}\right\}$
is included in the event $\{\widehat{w}_{\mathbf{X}}\notin\mathcal{S}_{n}\cap[-L_{n},L_{n}]^{p}\}$, we have
$\mathbb{P}(\widehat{w}_{\mathbf{X}}\in\mathcal{S}_{n})$ $\leq\mathbb{P}(\widehat{w}_{\mathbf{X}}\in\mathcal{S}_{n}\cap[-L_{n},L_{n}]^{p})+\mathbb{P}(A^{c})$
$\leq\mathbb{P}\left(\min_{\|w\|_{\infty}\leq L_{n}}H_{\ell}(w,\mathbf{X})\geq z+3\rho_{0}\right)+\mathbb{P}\left(\min_{w\in\mathcal{S}_{n}\cap[-L_{n},L_{n}]^{p}}H_{\ell}(w,\mathbf{X})\leq z+6\rho_{0}\right)+\varepsilon_{n}.$
For the first term, by Theorem 1, we obtain
$\mathbb{P}\left(\min_{\|w\|_{\infty}\leq L_{n}}H_{\ell}(w,\mathbf{X})\geq z+3\rho_{0}\right)$
$\leq$ $\mathbb{P}\left(\min_{\|w\|_{\infty}\leq L_{n}}H_{\ell}(w,\mathbf{W})\geq z+3\rho_{0}\right)+C_{1}(\max\{1,\rho_{0}^{-3}\})r_{R}(L_{n})$
$\leq$ $\mathbb{P}\left(\min_{w\in\mathbb{R}^{p}}H_{\ell}(w,\mathbf{W})\geq z+3\rho_{0}\right)+\varepsilon_{n}+C_{1}(\max\{1,\rho_{0}^{-3}\})r_{R}(L_{n})$
$\leq$ $2\varepsilon_{n}+C_{1}(\max\{1,\rho_{0}^{-3}\})r_{R}(L_{n}),$
where the last two inequalities hold by assumptions. For the second term, we have
$\mathbb{P}\left(\min_{w\in\mathcal{S}_{n}\cap[-L_{n},L_{n}]^{p}}H_{\ell}(w,\mathbf{X})\leq z+6\rho_{0}\right)\leq$ $\mathbb{P}\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{X})\leq z+6\rho_{0}\right)$
$\leq$ $\mathbb{P}\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{W})\leq z+18\rho_{0}\right)+C_{1}(\max\{1,\rho_{0}^{-3}\})r_{R}(L_{n})$
$\leq$ $\varepsilon_{n}+C_{1}(\max\{1,\rho_{0}^{-3}\})r_{R}(L_{n}),$
where the second last inequality follows by Theorem 1. The last inequality holds from the assumption. Therefore, the required result follows. ∎
A.3. Proof of Corollary 2 and Corollary 4
Proof.
Whichever corollary we prove, we verify that $r_{R}(L_{n})$ converges to zero with the setup. About the minimization in terms of $\delta\in(0,\omega_{n})$, we take $\delta=\delta_{n}\searrow 0$ as $n\to\infty$ such that $L_{n}/\delta_{n}\to\infty$ holds, which implies that $\log^{2/3}_{+}(L_{n}/\delta_{n})=o(\log n)$ and $\mathcal{M}_{R}(L_{n},\delta_{n})=o(1)$ with $\delta_{n}\searrow 0$ using the convexity of $R(\cdot)$. We can find such $\delta_{n}$, since $L_{n}$ has a logarithmic rate in $n$. We set $d=o(n^{\mu}/(\log n)^{d^{*}})$ with some $\mu,d^{*}\in\mathbb{R}$ and apply the definition of $\sigma_{n}$ in (7), then obtain
$\log^{2/3}_{+}(L_{n}/\delta_{n})L_{n}^{(\bar{\mathbf{q}}_{0}+3)/{3}}\sigma^{1/3}_{n}$
$=o(\log n)O(n^{(\bar{\mathbf{q}}_{0}+3)\alpha/{3}}(\log n)^{(\bar{\mathbf{q}}_{0}+3)\kappa/{3}})o(n^{2\mu/3}/(\log n)^{2d^{*}/3})n^{-1/6}$
$=o(n^{(\bar{\mathbf{q}}_{0}+3)\alpha/{3}+2\mu/3-1/6}(\log n)^{1-d^{*}/6+(\bar{\mathbf{q}}_{0}/3+1)\kappa}).$
Therefore, we obtain $\log^{2/3}_{+}(L_{n}/\delta_{n})L_{n}^{(\bar{\mathbf{q}}_{0}+3)/{3}}\sigma^{1/3}_{n}=o(1)$, when $\mu$ is no more than $21/4-(\bar{\mathbf{q}}_{0}+3)\alpha/$ 2 and $d^{*}$ is greater than or equal to $6+(2\bar{\mathbf{q}}_{0}+6)\kappa$. Using the result, we obtain
$r_{R}(L_{n})\lesssim\inf_{\rho\in(0,\bar{\rho})}\left\{\mathcal{M}_{\ell}(\rho)+\mathcal{D}^{3}_{\ell}(\rho)\cdot\zeta_{n}\right\},$
with some sequence $\zeta_{n}=o(1)$.
In the squared loss case, we obtain $\mathcal{M}_{\ell}(\rho)=0$ and $\mathcal{D}_{\ell}(\rho)\leq C$ with some $C>0$, hence we obtain the following as $n\to\infty$:
$r_{R}(L_{n})=\zeta_{n}=o(1).$ (14)
In the absolute loss and the Huber loss cases, we have $\mathcal{M}_{\ell}(\rho)\lesssim\rho$ and $\mathcal{D}_{\ell}(\rho)\lesssim\rho^{-2}$, which implies
$r_{R}(L_{n})\lesssim\inf_{\rho\in(0,\bar{\rho})}\left\{\rho+\rho^{-6}\cdot\zeta_{n}\right\}.$ (15)
Then, at the minimiation in $\rho$, we take $\rho=\rho_{n}=\zeta_{n}^{1/7}$, then we obtain
$r_{R}(L_{n})\lesssim\zeta_{n}^{1/7}=o(1).$ (16)
Finally, we obtain the statement. ∎
Appendix B Proof of Lemmas and Propositions
B.1. Preliminaries and Some Auxiliary Results
For the later use, we summarize the notations and the derivatives of $g(n^{-1}F_{\beta}(\cdot))$. Set $F_{\beta}(A)$ and $\bar{H}(w,A)$ as follows:
$F_{\beta}(A)$ $:=F_{\beta}((\bar{H}(w,A))_{w\in\mathcal{W}})=-\frac{1}{\beta}\log(\sum_{w\in\mathcal{W}}\exp(-\beta\bar{H}(w,A))),$
and
$\bar{H}(w,A)$ $:=n\times H(w,A)=\sum_{i=1}^{n}\ell_{i}((Aw)_{i})+n\cdot R(w).$
We introduce the notation $\langle\cdot\rangle_{H_{0}}$ as follows:
$\langle f\rangle_{H_{0}}:=\frac{\sum_{w\in\mathcal{W}}f(w,A)e^{-H_{0}(w,A)}}{\sum_{w\in\mathcal{W}}e^{-H_{0}(w,A)}}.$
Then, for any $g\in C^{3}(\mathbb{R})$,we have
$\partial_{i,j}g(n^{-1}F_{\beta}(A))=n^{-1}\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime}(n^{-1}F_{\beta}(A)),$
and
$\partial_{i,j}\partial_{i,k}g(n^{-1}F_{\beta}(A))=$ $n^{-2}\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime}(n^{-1}F_{\beta}(A))$
$+$ $n^{-1}g^{\prime}(n^{-1}F_{\beta}(A))\left(\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\right).$
where we define $\bar{H}_{\beta}:=\beta\cdot\bar{H}$ Also, we obtain the higher-order derivative as
$\partial_{i,j}\partial_{i,k}\partial_{i,\ell}g(n^{-1}F_{\beta}(A))$
$=m^{-3}\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime\prime}(n^{-1}F_{\beta}(A))$
$+$ $m^{-2}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime}(n^{-1}F_{\beta}(A))\left(\langle\partial_{i,j}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
$+$ $m^{-2}\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime}(n^{-1}F_{\beta}(A))\left(\langle\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,k}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
$+$ $m^{-2}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\cdot g^{\prime\prime}(n^{-1}F_{\beta}(A))\left(\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
$+$ $n^{-1}g^{\prime}(n^{-1}F_{\beta}(A))$
$\times$ $\left\{\langle\partial_{i,j}\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\partial_{i,k}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right.$
$-\beta\left(\langle\partial_{i,j}\partial_{i,\ell}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}+\langle\partial_{i,j}\bar{H}\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right.$
$\left.+\beta\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
$+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\left(\langle\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,k}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)$
$+\left.\beta\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\left(\langle\partial_{i,j}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}-\beta\langle\partial_{i,j}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}+\beta\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}\right)\right\}$
B.2. Proof of Lemma 5
Proof.
$W_{i}^{(B_{k})}$ is defined in the similar manner of $X_{i}^{(B_{k})}$. Let $(X_{i})^{(B_{k})}_{\ell}$ denote the $\ell$-th non-zero element of $X_{i}^{(B_{k})}$. We define
$\mathbf{Z}_{i,j}:=$ $(X_{1},\cdots,X_{i-1},(W_{i}+\sum_{\ell=1}^{j}(X_{i}^{(B_{\ell})}-W_{i}^{(B_{\ell})})),W_{i+1},\cdots,W_{n})^{\top},$ (17)
$\mathbf{Z}_{i,j}^{0}$ $:=\left(X_{1},\cdots,X_{i-1},(W_{i}-X_{i}^{(B_{j})}+\sum_{\ell=1}^{j}(X_{i}^{(B_{\ell})}-W_{i}^{(B_{\ell})})),W_{i+1},\cdots,W_{n}\right)^{\top}.$ (18)
Especially, when $j=0$ holds, $\mathbf{Z}_{i,0}$ becomes $(X_{1},\cdots,X_{i-1},W_{i},W_{i+1},\cdots,W_{n})^{\top}$.
By definition, we have
$E[f(\mathbf{X})]-E[f(\mathbf{W})]=\sum_{i=1}^{n}$ $\left(\sum_{j=1}^{k}E[(f(\mathbf{Z}_{i,j})-f(\mathbf{Z}^{0}_{i,j}))-(f(\mathbf{Z}_{i,j-1})-f(\mathbf{Z}^{0}_{i,j}))]\right).$
By third-order Taylor approximation, we have
$f(\mathbf{Z}_{i,j})$ $=f(\mathbf{Z}^{0}_{i,j})+(X^{(B_{j})}_{i})^{\top}\nabla_{i}f(\mathbf{Z}_{i,j}^{0})+\frac{1}{2}(X^{(B_{j})}_{i})^{\top}\nabla^{2}_{i}f(\mathbf{Z}_{i,j}^{0})X^{(B_{j})}_{i}$
$+\frac{1}{2}\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})dt$
and similarly,
$f(\mathbf{Z}_{i,j})$ $=f(\mathbf{Z}^{0}_{i,j})+(W^{(B_{j})}_{i})^{\top}\nabla_{i}f(\mathbf{Z}_{i,j}^{0})+\frac{1}{2}(W^{(B_{j})}_{i})^{\top}\nabla^{2}_{i}f(\mathbf{Z}_{i,j}^{0})W^{(B_{j})}_{i}$
$+\frac{1}{2}\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}W_{i,\ell}W_{i,\ell^{\prime}}W_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tW^{(B_{j})}_{i})dt$
where we define
$\nabla_{i}f(\mathbf{Z}^{0}_{i,j}):=\begin{bmatrix}\partial_{i,1}f(\mathbf{Z}^{0}_{i,j})\\ \vdots\\ \partial_{i,d}f(\mathbf{Z}^{0}_{i,j})\\ \end{bmatrix},\quad\nabla^{2}_{i}f(\mathbf{Z}^{0}_{i,j}):=\begin{bmatrix}\partial^{2}_{i,1}f(\mathbf{Z}^{0}_{i,j})&\cdots&\partial_{i,1}\partial_{i,d}f(\mathbf{Z}^{0}_{i,j})\\ \vdots&\ddots&\vdots\\ \partial_{i,1}\partial_{i,d}f(\mathbf{Z}^{0}_{i,j})&\cdots&\partial^{2}_{i,d}f(\mathbf{Z}^{0}_{i,j})\\ \end{bmatrix}.$
Therefore, we have
$E[f(\mathbf{X})]-E[f(\mathbf{W})]$
$=$ $\sum_{i=1}^{n}\left(\sum_{j=1}^{k}E[(f(\mathbf{Z}_{i,j})-f(\mathbf{Z}^{0}_{i,j}))-(f(\mathbf{Z}_{i,j-1})-f(\mathbf{Z}^{0}_{i,j}))]\right)$
$=$ $\sum_{i=1}^{n}\sum_{j=1}^{k}\sum_{\ell=1}^{|B_{j}|}E[((X^{(B_{j})}_{i})_{\ell}-(W^{(B_{j})}_{i})_{\ell})\partial_{i,(B_{j})_{\ell}}f(\mathbf{Z}^{0}_{i,j})]$
$+$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{k}\sum_{\ell=1}^{|B_{j}|}\sum_{\ell^{\prime}=1}^{|B_{j}|}E[((X^{(B_{j})}_{i})_{\ell}(X^{(B_{j})}_{i})_{\ell^{\prime}}-(W^{(B_{j})}_{i})_{\ell}(W^{(B_{j})}_{i})_{\ell^{\prime}})\partial_{i,(B_{j})_{\ell}}\partial_{,i(B_{j})_{\ell^{\prime}}}f(\mathbf{Z}^{0}_{i,j})]$
$+$ $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{k}\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})dt$
-- $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{k}\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}W_{i,\ell}W_{i,\ell^{\prime}}W_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}f(\mathbf{Z}_{i,j}^{0}+tW^{(B_{j})}_{i})dt$
By the assumption on block dependence, for any $j$ and $\ell\in[|B_{j}|]$, we have
$E[((X^{(B_{j})}_{i})_{\ell}-(W^{(B_{j})}_{i})_{\ell})\partial_{i,(B_{j})_{\ell}}f(\mathbf{Z}^{0}_{i,j})]$ $=E[E[((X^{(B_{j})}_{i})_{\ell}-(W^{(B_{j})}_{i})_{\ell})|\mathbf{Z}^{0}_{i,j}]\partial_{i,(B_{j})_{\ell}}f(\mathbf{Z}^{0}_{i,j})]$
$=E[E[((X^{(B_{j})}_{i})_{\ell}-(W^{(B_{j})}_{i})_{\ell})]\partial_{i,(B_{j})_{\ell}}f(\mathbf{Z}^{0}_{i,j})]$
$=0.$
In the same way, or any $j$ and $\ell\in[B_{j}]$, we have
$E[((X^{(B_{j})}_{i})_{\ell}(X^{(B_{j})}_{i})_{\ell^{\prime}}-(W^{(B_{j})}_{i})_{\ell}(W^{(B_{j})}_{i})_{\ell^{\prime}})\partial_{i,(B_{j})_{\ell}}\partial_{i,(B_{j})_{\ell^{\prime}}}f(\mathbf{Z}^{0}_{i,j})]=0.$
Therefore, we have the required result. ∎
B.3. Proof of Proposition 7
To derive the result of , we introduce two lemmas. First, we establish that the difference in the expectation of certain functions can be upper bounded by moments of their inputs.
Lemma 13.
Consider two sequences of i.i.d. $\mathbb{R}^{p}$-valued random vectors $X_{1},...,X_{n}$ and $W_{1},...,W_{n}$. Suppose that each $\{X_{i}\}_{i=1}^{n}$ and $\{W_{i}\}_{i=1}^{n}$ is $d$-block dependent, and $E[X_{i}]=E[W_{i}]$ and $E[X_{i}X_{i}^{\top}]=E[W_{i}W_{i}^{\top}]$ hold. Let $\mathbf{X}$ and $\mathbf{W}$ denote matrices $(X^{\top}_{1},\cdots,X^{\top}_{n})^{\top}$ and $(W^{\top}_{1},\cdots,W^{\top}_{n})^{\top}$, respectively. Then, there exists some $C_{0}=C_{0}(\tau,q)>0$ such that for any finite set $\mathcal{W}\subset[-L_{n},L_{n}]^{p}$ with $L_{n}\geq 1$, any $\rho\in(0,1)$, and any $g\in C^{3}(\mathbb{R}^{n\times d})$ satisfying $\max_{i\in[n],j\in[p]}|E[\partial_{i,j}g(\mathbf{Z}^{0}_{i})]|<\infty$ and $\max_{i\in[n],j,k\in[p]}|E[\partial_{i,j}\partial_{i,k}g(\mathbf{Z}^{0}_{i})]|<\infty$, we have
$|E[g(n^{-1}F_{\beta}((\bar{H}_{\ell_{\rho}}(w,\mathbf{X})))_{w\in\mathcal{W}}]-E[g(n^{-1}F_{\beta}((\bar{H}_{\ell_{\rho}}(w,\mathbf{W})))_{w\in\mathcal{W}}]|$
$\leq C_{0}(\tau,q)K_{g}\max\{1,\beta^{2}\}L_{n}^{\bar{\mathbf{q}}_{0}+3}D^{3}_{\ell}(\rho)$
×$\times\left(\frac{1}{n}\sum_{i=1}^{n}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]\right.$
$+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]$
$+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{\bar{\mathbf{q}}_{0}+3}\right]$
$\left.+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]\right).$
where we set $\tau\leq\frac{n}{d}\leq\frac{1}{\tau}$ and denote $C^{j}_{i\ell}$ as the $(i,\ell)$ element of $\mathbf{Z}_{i,j}^{0}$. The summation of the last term expresses the sum of all the combinations of $\mathbf{{\bar{q}}}_{0}$ random variables exerted from the $i$-th row of $\mathbf{Z}_{i,j}^{0}$ except those in $B_{j}$.
Second, we analyze and derive the upper bound of the complicated summation of the moment in Lemma 13.
Lemma 14.
Suppose Assumptions 1 and 2 hold. Let $C_{i}^{j}$ denote the $i$-th row of $\mathbf{Z}^{0}_{ij}$. Consider two sequences of i.i.d. $\mathbb{R}^{p}$-valued random vectors $X_{1},...,X_{n}$ and $W_{1},...,W_{n}$. Suppose that each $\{X_{i}\}_{i=1}^{n}$ and $\{W_{i}\}_{i=1}^{n}$ is $d$-block dependent. Suppose $\bar{\mathbf{q}}_{0}\leq p$ holds. Then, for any $B_{j}$, we have
$\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]$
$\lesssim_{\tau,q}\frac{d^{(\bar{\mathbf{q}}_{0}/2)+1}}{n}\max\{\max_{1\leq k\leq p}(\|X_{i,k}\|_{2^{(\bar{\mathbf{q}}_{0}/2)}})^{(\bar{\mathbf{q}}_{0})},\max_{1\leq k\leq p}(\|W_{i,k}\|_{2^{(\bar{\mathbf{q}}_{0}/2)}})^{(\bar{\mathbf{q}}_{0})}\}+(\max_{1\leq i\leq k}\lambda_{\max}(\Sigma_{|B_{i}|}))^{(\bar{\mathbf{q}}_{0}/2)},$
where the summation expresses the sum of all the combinations of $\mathbf{{\bar{q}}}_{0}$ random variables with repetition, exerted from the $i$-th row of $\mathbf{Z}_{i,j}^{0}$ except those in $B_{j}$.
Now, we are ready to prove Proposition 7.
Proof of Proposition 7.
From Lemma 13 and 14, we have
$|E[g(n^{-1}F_{\beta}((\bar{H}_{\ell_{\rho}}(w,\mathbf{X})))_{w\in\mathcal{W}}]-E[g(n^{-1}F_{\beta}((\bar{H}_{\ell_{\rho}}(w,\mathbf{W})))_{w\in\mathcal{W}}]|$
$\lesssim_{\tau,q,M}K_{g}\max\{1,\beta^{2}\}L_{n}^{\bar{\mathbf{q}}_{0}+3}D^{3}_{\ell}(\rho)$
×$\times\left\{\frac{1}{n}\sum_{i=1}^{n}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]\right.$
$+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]$
$+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{\bar{\mathbf{q}}_{0}+3}\right]$
$+\left.\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C_{ik_{1}}\cdots C_{ik_{\bar{\mathbf{q}}_{0}}}\right]\right\}.$
Then, we need to derive the upper bounds for the four terms.
By definition, for each $j\in\{1,\cdots,k\}$, we have
$E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(X^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]=$ $E\left[\sum_{\ell=1}^{|B_{j}|}\sum_{\ell^{\prime}=1}^{|B_{j}|}\sum_{\ell^{\prime\prime}=1}^{|B_{j}|}|(X^{(B_{j})}_{i})_{\ell}(X^{(B_{j})}_{i})_{\ell^{\prime}}(X^{(B_{j})}_{i})_{\ell^{\prime\prime}}|\right]$
$\leq$ $|B_{j}|^{3}\max_{\ell,\ell^{\prime},\ell^{\prime\prime}\in[|B_{j}|]}E\left[|(X^{(B_{j})}_{i})_{\ell}(X^{(B_{j})}_{i})_{\ell^{\prime}}(X^{(B_{j})}_{i})_{\ell^{\prime\prime}}|\right]$
By the Cauchy-Schwartz inequality, for any ,,′′$\ell,\ell^{\prime},\ell^{\prime\prime}\in[|B_{j}|]$, we have
$E\left[|(X^{(B_{j})}_{i})_{\ell}(X^{(B_{j})}_{i})_{\ell^{\prime}}(X^{(B_{j})}_{i})_{\ell^{\prime\prime}}|\right]$ $\leq\|(X^{(B_{j})}_{i})_{\ell}(X^{(B_{j})}_{i})_{\ell^{\prime}}\|_{2}\|(X^{(B_{j})}_{i})_{\ell^{\prime\prime}}\|_{2}$
$\leq\|(X^{(B_{j})}_{i})_{\ell}\|_{4}\|(X^{(B_{j})}_{i})_{\ell^{\prime}}\|_{4}\|(X^{(B_{j})}_{i})_{\ell^{\prime\prime}}\|_{4}$
$\leq\max_{\ell\in[|B_{j}|]}\|(X^{(B_{j})}_{i})_{\ell}\|^{3}_{4}$
Similarly, for each $j\in\{1,\cdots,k\}$, we also have
$E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(X^{(B_{j})}_{i})_{\ell}\right|\right)^{\bar{\mathbf{q}}_{0}+3}\right]\leq|B_{j}|^{\bar{\mathbf{q}}_{0}+3}\max_{\ell\in[|B_{j}|]}\|(X^{(B_{j})}_{i})_{\ell}\|^{\bar{\mathbf{q}}_{0}+3}_{2^{(\bar{\mathbf{q}}_{0}+4)/2}}.$
Then, from Lemma 6 and 14, we obtain
$\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]$ $\lesssim\frac{pd^{2}}{n^{3/2}}\max\left\{\max_{1\leq k\leq p}\|X_{i,k}\|^{3}_{4},\max_{1\leq k\leq p}\|W_{i,k}\|^{3}_{4}\right\},$
and
$\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{\bar{\mathbf{q}}_{0}+3}\right]$
$\lesssim\frac{pd^{\bar{\mathbf{q}}_{0}+2}}{n^{(\bar{\mathbf{q}}_{0}+3)/2}}\max\left\{\max_{1\leq k\leq p}\|X_{i,k}\|^{\bar{\mathbf{q}}_{0}+3}_{2^{(\bar{\mathbf{q}}_{0}+4)/2}},\max_{1\leq k\leq p}\|W_{i,k}\|^{\bar{\mathbf{q}}_{0}+3}_{2^{(\bar{\mathbf{q}}_{0}+4)/2}}\right\},$
Furthermore, we have
$\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C_{ik_{1}}\cdots C_{ik_{\bar{\mathbf{q}}_{0}}}\right]$
$\lesssim_{\tau,q}$ $\frac{pd^{3+(\bar{\mathbf{q}}_{0}/2)}}{n^{5/2}}\max\left\{\max_{1\leq k\leq p}\|X_{i,k}\|^{3}_{4},\max_{1\leq k\leq p}\|W_{i,k}\|^{3}_{4}\right\}\left(\max\{\max_{1\leq k\leq p}\|X_{i,k}\|^{(\bar{\mathbf{q}}_{0})}_{2^{(\bar{\mathbf{q}}_{0}/2)}},\max_{1\leq k\leq p}\|W_{i,k}\|^{(\bar{\mathbf{q}}_{0})}_{2^{(\bar{\mathbf{q}}_{0}/2)}}\}\right)$
$+$ $\frac{pd^{2}}{n^{3/2}}\max\left\{\max_{1\leq k\leq p}\|X_{i,k}\|^{3}_{4},\max_{1\leq k\leq p}\|W_{i,k}\|^{3}_{4}\right\}\left(\max_{1\leq i\leq k}\lambda_{\max}(\Sigma_{|B_{i}|})\right)^{\bar{\mathbf{q}}_{0}/2}.$
Therefore, the required result holds. ∎
B.4. Proof of Lemma 8 and Lemma 9
Proof of Lemma 8.
We construct a finite set $\mathcal{S}_{n,\delta}$ such that $|\mathcal{S}_{n,\delta}|\leq\lceil 2L_{n}/\delta\rceil^{p}$. Fix $\delta\in(0,1)$. Define $\delta\mathbf{Z}$ as the set of integer multiples of $\delta$, i.e., $\delta\mathbf{Z}:=\{\dots,-2\delta,-\delta,0,\delta,2\delta,\dots\}$. We then define $\mathcal{S}_{n,\delta}$ as the intersection of $\delta\mathbf{Z}$ with $\mathcal{S}_{n}$. By construction, the number of points in each dimension of $\mathcal{S}_{n,\delta}$ is bounded above by $\lceil 2L_{n}/\delta\rceil$, implying that $|\mathcal{S}_{n,\delta}|\leq\lceil 2L_{n}/\delta\rceil^{p}$. Let $w_{*}$ and $w_{*,\delta}$ denote the arguments of the minimum in $\mathcal{S}_{n}$ and $\mathcal{S}_{n,\delta}$ for $H_{\ell_{\rho}}(w,S)$, respectively. By the definition of $\mathcal{S}_{n,\delta}$, it follows that $\|w_{*}-w_{*,\delta}\|\leq\delta$. Then, we obtain
$\left|\mathbb{E}\left[g\left(\min_{w\in\mathcal{S}_{n,\delta}}H_{\ell_{\rho}}(w,\mathbf{X})\right)\right]-\mathbb{E}\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell_{\rho}}(w,\mathbf{X})\right)\right]\right|$
$\leq\|g^{\prime}\|_{\infty}\mathbb{E}\left[\left(H_{\ell_{\rho}}(w_{*,\delta},\mathbf{X})-H_{\ell_{\rho}}(w_{*},\mathbf{X})\right)_{+}\right]$
$\leq\|g^{\prime}\|_{\infty}\left(\frac{1}{n}\mathbb{E}\left[\left|\sum_{i=1}^{n}\left(\ell_{i;\rho}((\mathbf{X}w_{*,\delta})_{i})-\ell_{i;\rho}((\mathbf{X}w_{*})_{i})\right)\right|\right]+\sup_{\begin{subarray}{c}w,w^{\prime}\in[-L_{n},L_{n}]^{p}\\ \|w-w^{\prime}\|_{\infty}\leq\delta\end{subarray}}|R(w)-R(w^{\prime})|\right).$
From Assumption 2, we have
$|\ell_{i;\rho}((\mathbf{X}w_{*,\delta})_{i})-\ell_{i;\rho}((\mathbf{X}w_{*})_{i})|\leq 2\mathcal{D}_{\ell}(\rho)\left|(\mathbf{X}(w_{*}-w_{*,\delta}))_{i}\right|L_{\ell_{i}}^{q_{1}}\left(1+|(\mathbf{X}w_{*})_{i}|^{q_{0}q_{1}}+|(\mathbf{X}w_{*,\delta})_{i}|^{q_{0}q_{1}}\right).$
Thus, the inequality yields
$\frac{1}{n}E\left[\left|\sum_{i=1}^{n}(\ell_{i;\rho}((\mathbf{X}w_{*,\delta})_{i})-\ell_{i;\rho}((\mathbf{X}w_{*})_{i}))\right|\right]$
$\leq 2\mathcal{D}_{\ell}(\rho)E\left[\|\mathbf{X}(w_{*}-w_{*,\delta})\|_{\infty}(1+\|\mathbf{X}w_{*}\|_{\infty}^{q_{0}q_{1}}+\|\mathbf{X}w_{*,\delta}\|_{\infty}^{q_{0}q_{1}})\frac{1}{n}\sum_{i=1}^{n}L^{q_{1}}_{\ell_{i}}\right]$
$\leq 2\mathcal{D}_{\ell}(\rho)\delta E\left[\left(\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|\right)(1+\|\mathbf{X}w_{*}\|_{\infty}^{q_{0}q_{1}}+\|\mathbf{X}w_{*,\delta}\|_{\infty}^{q_{0}q_{1}})\frac{1}{n}\sum_{i=1}^{n}L^{q_{1}}_{\ell_{i}}\right]$
$\leq 2\mathcal{D}_{\ell}(\rho)\delta L^{q_{0}q_{1}}_{n}Av(\{L^{q_{1}}_{\ell_{i}}\})\left(E\left[\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|\right]+E\left[\left(\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|\right)^{q_{0}q_{1}+1}\right]\right)$
where the second inequality follows from $\|w_{*}-w_{*,\delta}\|\leq\delta$ and the third inequality follows from the $\ell_{\infty}$ bound of the parameter space $\mathcal{S}_{n}$. Therefore, the required result holds.
Proof of Lemma 9.
It simply follows from Proposition 4.1 in Han and Shen (2023). ∎
Appendix C Proof for Applications
C.1. Proof of Lemma 10
Proof.
Let $\widehat{w}^{(s)}$ denote an the column leave-one-out estimator, i.e., $\widehat{w}^{(s)}:=\arg\min_{w\in\mathbf{R}^{p}:w_{s}=0}H(w)$. Without assuming Assumption iii., as in the proof of Proposition 3.11 in Han and Shen (2023), we obtain
$|\widehat{w}_{\mathbf{X},s}|\lesssim\lambda^{-1}\left|\sum_{i=1}^{n}X_{i,s}\psi^{\prime}_{0}(X^{\top}_{i,-s}\widehat{w}^{(s)}_{\mathbf{X},-s}-\xi_{i})\right|+\|\theta_{0}\|_{\infty}$
where we define $\widehat{w}_{\mathbf{X},s}$ and $\widehat{w}^{(s)}_{\mathbf{X},-s}$ as the $s-$th element of $\widehat{w}_{\mathbf{X}}$ and the one removing the $s-$th element of $\widehat{w}^{(s)}$ from the original vector, respectively. Therefore, for $k\geq 2$, we obtain
$E[|\widehat{w}_{\mathbf{X},s}|^{k}]\lesssim_{k}$ $\lambda^{-k}E\left[\left|\sum_{i=1}^{n}X_{i,s}\psi^{\prime}_{0}(X^{\top}_{i,-s}\widehat{w}^{(s)}_{\mathbf{X},-s}-\xi_{i})\right|^{k}\right]+\|\theta_{0}\|^{k}_{\infty}$
[…]
where $\{\varepsilon_{i}\}$ are i.i.d. Rademacher random variables independent of $\mathbf{X}$. Inequality (i) follows from Assumption 4-i.. We applied the symmetrization inequality of Theorem 3.1.21 in Giné and Nickl (2016) and Khintchine’s inequality to inequality (ii) and inequality (iii), respectively. From the Jensen’s inequality, we obtain
$E\left[\left|\sum_{i=1}^{n}X^{2}_{i,s}\right|^{k/2}\right]\leq\left(E\left[\left|\sum_{i=1}^{n}X^{2}_{i,s}\right|^{\lceil k/2\rceil}\right]\right)^{k/2{\lceil k/2\rceil}}.$
By frequently applying the Lyapunov inequality, we have
$E\left[\left|\sum_{i=1}^{n}X^{2}_{i,s}\right|^{\lceil k/2\rceil}\right]\leq E[|X^{0}_{i,s}|^{2\lceil k/2\rceil}].$
Therefore, the required result holds. ∎
C.2. Proof of Theorem 11
Proof.
It holds from Lemma 10 and Markov inequality that
$\mathbb{P}(\|\widehat{w}^{R}_{\mathbf{X},j}\|_{\infty}>L_{n})\leq L_{n}^{-k^{*}}E\max_{j\in[n]}|\widehat{w}^{R}_{\mathbf{X},j}|^{k^{*}}\lesssim_{L_{0},pk^{*}\lambda}(pL^{-k^{*}}_{n})((M_{k^{*};\mathbf{X}})^{k/2\lceil k/2\rceil}+\|\theta_{0}\|^{k^{*}}_{\infty})$
As in Han and Shen (2023), we apply Theorem 3. In the robust case, $L_{\ell_{i}}=CL_{0}(1+\xi_{i})$, $\mathbf{q}=0$ , $\mathcal{D}_{\ell}(\rho)=CL_{0}/\rho^{2}$, $\mathcal{M}_{\ell}(\rho)=CL_{0}\rho$ and $\bar{\rho}=1/(CL_{0})$ hold from Lemma B.1 in Han and Shen (2023). Then, by setting $\rho$ and $\delta$ appropriately, we have
$r_{R}(L_{n})\lesssim_{L_{0}}\left(L_{n}\log^{2/3}(n)\sigma^{1/3}_{n}\right)^{1/7}$
Therefore, there exists $K>0$ depending on $k^{*},M_{k^{*};A},\tau,\lambda$ such that
$\mathbb{P}(\widehat{w}^{R}_{\mathbf{X},j}\in\mathcal{S}_{n})\leq 4\varepsilon_{n}+K\left\{(nL^{-k^{*}}_{n})(1\vee\|\theta_{0}\|^{k^{*}})+(1\vee\rho^{-3}_{0})\left(L_{n}\log^{2/3}(nL_{n})\sigma^{1/3}_{n}\right)^{1/7}\right\}$
Now setting $L_{n}:=(n/d^{4})^{1/6-\delta}$ where we define $\delta:=\frac{1}{6}\left(\frac{7\varepsilon k^{*}-42}{7\varepsilon k^{*}+1}\right)$. Therefore, we obtain
$\mathbb{P}(\widehat{w}^{R}_{\mathbf{X},j}\in\mathcal{S}_{n})\leq 4\varepsilon_{n}+K(1+\|\theta_{0}\|^{k^{*}}_{\infty}+\rho^{-3}_{0})n^{-\frac{1}{45}}.$
C.3. Proof of Theorem 12
Proof.
All the conditions for the Theorem 4.1 in Thrampoulidis et al. (2018) trivially hold from Theorem 11 and assumptions of Theorem 12 as in Han and Shen (2023). ∎
Appendix D Proof of Auxiliary Results
D.1. Proof of Lemma 13
Proof.
Define $\bar{H}_{\beta}:=\beta\cdot\bar{H}$ and $F_{\beta}(A):=F_{\beta}(\bar{H}(w,A)_{w\in\mathcal{W}})$. As an initial step, we apply the result from Lemma 5, reducing the difference to the third derivatives of the functions involved. As shown in Appendix B.1, the third derivative of the function $g(n^{-1}F_{\beta}(\cdot))$ can be bounded by the derivatives of the loss function $\bar{H}$. For any $i\in\{1,\cdots,n\}$ and $j,k,\ell\in\{1,\cdots,d\}$, we have the following upper bound of derivatives of $\bar{H}$.
$|\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq L_{n}D_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{1}}\rangle_{\bar{H}_{\beta}})$
$|\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq L^{2}_{n}D_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{2}}\rangle_{\bar{H}_{\beta}})$
$|\langle\partial_{i,j}\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq L^{3}_{n}D_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{3}}\rangle_{\bar{H}_{\beta}})$
where we define $\bar{H}^{(i)}(w,A):=\ell_{i:\rho}((Aw)_{i})\geq 0$. so $\bar{H}(w,A)=\sum_{i=1}^{n}\bar{H}^{(i)}(w,A)+nR(w)$. Therefore, any $i\in\{1,\cdots,n\}$ and $j,k,\ell\in\{1,\cdots,d\}$, we have
$|\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq L^{3}_{n}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{1}}\rangle_{\bar{H}_{\beta}})^{3}$
$\leq 2^{2}L^{3}_{n}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{3q_{1}}\rangle_{\bar{H}_{\beta}})$
$|\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,k}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq\langle|w_{j}\partial\ell_{i:\rho}((Aw)_{i})|\rangle_{\bar{H}_{\beta}}\langle|w_{k}||w_{\ell}||\partial\ell_{i:\rho}((Aw)_{i})|^{2}\rangle_{\bar{H}_{\beta}}$
$\leq L^{3}_{n}\langle|\partial\ell_{i:\rho}((Aw)_{i})|^{3}\rangle_{\bar{H}_{\beta}}$
$\leq 2^{2}L^{3}_{n}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{3q_{1}}\rangle_{\bar{H}_{\beta}})$
$|\langle\partial_{i,j}\bar{H}\partial_{i,k}\bar{H}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq L^{3}_{n}\langle|\partial\ell_{i:\rho}((Aw)_{i})|^{3}\rangle_{\bar{H}_{\beta}}$
$\leq 2^{2}L^{3}_{n}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{3q_{1}}\rangle_{\bar{H}_{\beta}})$
$|\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}\langle\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq\langle|w_{j}w_{k}\partial^{2}\ell_{i:\rho}((Aw)_{i})|\rangle_{\bar{H}_{\beta}}\langle|w_{\ell}\partial\ell_{i:\rho}((Aw)_{i})|\rangle_{\bar{H}_{\beta}}$
$\leq L_{n}^{3}\langle|\partial^{2}\ell_{i:\rho}((Aw)_{i})|\rangle_{\bar{H}_{\beta}}\langle|\partial\ell_{i:\rho}((Aw)_{i})|\rangle_{\bar{H}_{\beta}}$
$\leq L_{n}^{3}D^{2}_{\ell}(\rho)\langle 1+|\ell_{i:\rho}((Aw)_{i})|^{q_{2}}\rangle_{\bar{H}_{\beta}}\langle 1+|\ell_{i:\rho}((Aw)_{i})|^{q_{1}}\rangle_{\bar{H}_{\beta}}$
$\lesssim L_{n}^{3}D^{2}_{\ell}(\rho)\langle 1+|\ell_{i:\rho}((Aw)_{i})|^{q_{1}+q_{2}}\}\rangle_{\bar{H}_{\beta}}$
$\lesssim L_{n}^{3}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{1}+q_{2}}\rangle_{\bar{H}_{\beta}})$
$|\langle\partial_{i,j}\partial_{i,\ell}\bar{H}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq|w_{j}w_{\ell}w_{k}|\langle|\partial\ell^{2}_{i:\rho}((Aw)_{i})||\partial\ell_{i:\rho}((Aw)_{i})|\rangle_{\bar{H}_{\beta}}$
$\leq L^{3}_{n}D^{2}_{\ell}(\rho)\langle(1+|\ell_{i:\rho}((Aw)_{i})|^{q_{2}})(1+|\ell_{i:\rho}((Aw)_{i})|^{q_{1}})\rangle_{\bar{H}_{\beta}}$
$\lesssim L_{n}^{3}D^{2}_{\ell}(\rho)\langle 1+|\ell_{i:\rho}((Aw)_{i})|^{q_{1}+q_{2}}\}\rangle_{\bar{H}_{\beta}}$
$\lesssim L_{n}^{3}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{1}+q_{2}}\rangle_{\bar{H}_{\beta}})$
$|\langle\partial_{i,j}\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq|w_{j}||w_{k}||w_{\ell}|\langle|\partial^{3}\ell_{i:\rho}((Aw)_{i})|\rangle_{\bar{H}_{\beta}}$
$\leq L_{n}^{3}D_{\ell}(\rho)\langle 1+|\ell_{i:\rho}((Aw)_{i})|^{q_{3}}\rangle_{\bar{H}_{\beta}}$
$\leq L_{n}^{3}D_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{3}}\rangle_{\bar{H}_{\beta}}$
Hence, for $\beta\geq 1$, we have the following upper bound for the $\partial_{i,j}\partial_{i,k}\partial_{i,\ell}g(n^{-1}F_{\beta}(A))$:
$|\partial_{i,j}\partial_{i,k}\partial_{i,\ell}g(n^{-1}F_{\beta}(A))|\lesssim K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)\left[1+\langle\bar{H}^{(i)}(w,A)^{\mathbf{q}}\rangle_{\bar{H}_{\beta}}\right]$
where we define $\mathbf{q}=\max\{3q_{1},2q_{1}+2q_{2},q_{3}\}$. So, we obtain
$E[|\partial_{i,(B_{j})_{\ell}}\partial_{i,(B_{j})_{\ell^{\prime}}}\partial_{i,(B_{j})_{\ell^{\prime\prime}}}g(n^{-1}F_{\beta}(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))||X^{(B_{j})}_{i}]$
$\lesssim$ $K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)\left[1+E[\langle(\bar{H}^{(i)}(w,\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta}}|X^{(B_{j})}_{i}]\right]$
By using Chebysev’s association inequality, we have
$E[\langle(\bar{H}^{(i)}(w,\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta}}|X^{(B_{j})}_{i}]\leq E[\langle(\bar{H}^{(i)}(w,\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}]$
where we define $H_{\beta;-i}:=\beta(\bar{H}(w,C)-\bar{H}^{(i)}(w,C))$ and $C$ is a $n\times d$ matrix . Therefore, we have
$\begin{split}&E[|\partial_{i,(B_{j})_{\ell}}\partial_{i,(B_{j})_{\ell^{\prime}}}\partial_{i,(B_{j})_{\ell^{\prime\prime}}}g(n^{-1}F_{\beta}(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))||X^{(B_{j})}_{i}]\\ \lesssim&K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)\left[1+E[\langle(\ell_{i;\rho}(((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}]\right]\end{split}$ (19)
In the second step, we derive an upper bound for the moments of the random variables by leveraging the properties of the loss function. Assumption 2 (1) and (2) yield
$\begin{split}&E[\langle(\ell_{i;\rho}(((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}]\\ &\leq E[\langle(\|\ell_{i;\rho}(((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i})-\ell_{i}(((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i})\|_{\infty}\\ &\quad+\ell_{i}(((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}]\\ &\lesssim_{q}(1+E[\langle(\ell_{i}(((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}])\\ &\lesssim_{q}(1+E[\langle L^{\mathbf{q}}_{\ell_{i}}(1+|((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i}|^{\mathbf{q}_{0}})\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}])\\ &\lesssim_{q}L^{\mathbf{q}}_{\ell_{i}}\left(1+E\left[\langle|((\mathbf{Z}_{i,j}^{0})^{\top}w)_{i}|^{\mathbf{q}_{0}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}\right]+E\left[\langle|t(X^{(B_{j})}_{i})^{\top}w|^{\mathbf{q}_{0}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}\right]\right)\end{split}$ (20)
where we define $\mathbf{q}_{0}=q_{0}\mathbf{q}$ and $\bar{\mathbf{q}}_{0}:=2\lceil\mathbf{q}_{0}/2\rceil$, respectively. By trivial calculation, we have
$|t(X^{(B_{j})}_{i})^{\top}w|^{\mathbf{q}_{0}}\leq t^{\mathbf{q}_{0}}L^{\mathbf{q}_{0}}_{n}\left(\sum_{m=1}^{|B_{j}|}|(X_{i}^{(B_{j})})_{m}|\right)^{\mathbf{q}_{0}}.$ (21)
Moreover, let $C^{j}_{i}$ denote the $i$-th row of $\mathbf{Z}_{i,j}^{0}$. Then, we have
$\begin{split}E\left[\langle|((\mathbf{Z}_{i,j}^{0})^{\top}w)_{i}|^{\bar{\mathbf{q}}_{0}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}\right]\leq&E\left[\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\mathbf{q}_{0}}}\langle w_{k_{1}}\cdots w_{k_{\bar{\mathbf{q}}_{0}}}\rangle_{\bar{H}_{\beta;-i}}\right]\\ \leq&\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]E\left[\langle w_{k_{1}}\cdots w_{k_{\bar{\mathbf{q}}_{0}}}\rangle_{\bar{H}_{\beta;-i}}\right]\\ \leq&L_{n}^{\bar{\mathbf{q}}_{0}}\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]\end{split}$ (22)
where the second inequality follows from i.i.d. property of $X_{i}$ and $W_{i}$. Therefore, from (19) to (22), we have
$\begin{split}&E[|\partial_{i,(B_{j})_{\ell}}\partial_{i,(B_{j})_{\ell^{\prime}}}\partial_{i,(B_{j})_{\ell^{\prime\prime}}}g(n^{-1}F_{\beta}(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))|X^{(B_{j})}_{i}]\\ \lesssim&K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)\left[1+L^{\mathbf{q}}_{\ell_{i}}L_{n}^{\bar{\mathbf{q}}_{0}}\left(1+\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]+t^{\bar{\mathbf{q}}_{0}}\left(\sum_{m=1}^{|B_{j}|}|(X^{(B_{j})}_{i})_{m}|\right)^{\bar{\mathbf{q}}_{0}}\right)\right]\end{split}$ (23)
holds. The second term expresses the sum of all combinations of $\mathbf{{\bar{q}}}_{0}$ random variables exerted from the $i$-th row of $\mathbf{Z}_{i,j}^{0}$ except those in $B_{j}$. Combining Lemma 5 with (23), we attain
$\left|E\left[\int^{1}_{0}(1-t)^{2}\sum_{\ell=1}^{|B_{j}|}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}g(n^{-1}F_{\beta}(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))dt\right]\right|$
[…]
$\lesssim$ $K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(X^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]$
$+$ $K_{g}n^{-1}\beta^{2}L_{n}^{\bar{\mathbf{q}}_{0}+3}D^{3}_{\ell}(\rho)L^{\mathbf{q}}_{\ell_{i}}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(X^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\left(1+\left(\sum_{\ell=1}^{|B_{j}|}\left|(X^{(B_{j})}_{i})_{\ell}\right|\right)^{\bar{\mathbf{q}}_{0}}+\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]\right)\right].$
Therefore, the required result holds. ∎
D.2. Proof of Lemma 14
Proof.
Let $C^{j}_{i}$ denote the $i$-th row of $\mathbf{Z}_{i,j}^{0}$. By definition, there exists $\alpha\in\mathbb{N}$ such that $\bar{\mathbf{q}}_{0}=2\alpha$. Throughout this proof, we assume $\bar{\mathbf{q}}_{0}\leq p$. The summation, $\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}$, expresses the sum of all combinations of random variables $\mathbf{{\bar{q}}}_{0}$ with repetition, exerted from the $i$-th row of $\mathbf{Z}_{i,j}^{0}$ except those in $B_{j}$. Let $\ell$ denote the number of different dependent index cells in which random variables are.
We divide this summation into three parts based on $\ell$, i.e. (i)$11\leq\ell\leq\bar{\mathbf{q}}_{0}-$ 1, (ii) $\ell=\bar{\mathbf{q}}_{0}-\alpha$ and (iii) $(\bar{\mathbf{q}}_{0}+1)-\alpha\leq\ell\leq\bar{\mathbf{q}}_{0}$, and denote corresponding summation as $A_{1}$, $A_{2}$ and $A_{3}$, respectively.
$\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]=A_{1}+A_{2}+A_{3}.$
We consider three cases above and derive upper bounds for them.
  1. (i)
    First, we consider the number of terms included in this case. After deriving the uniform upper bound of expectations, we can upper bound $A_{1}$ by the multiplication of the number of terms and the uniform bound.
    Let $N_{1}$ denote the number of terms included in this case. To derive the number of terms included in this cell, we consider the following steps:
    1. Step 1
      Fix $\ell$ and pick up a specific $k_{1}$ from $[p]$ except elements in $B_{j}$. The possible ways to select $C^{j}_{ik_{1}}$ is $p-|B_{j}|$. Let $B^{*}_{1,\ell}$ denote the dependent index cell of $C^{j}_{ik_{1}}$.
    2. Step 2
      Set $\gamma_{1,\ell}$ as the number of variables included in the same dependent index cell as $C^{j}_{ik_{1}}$. Because the maximum cardinality of dependent index cells is $d$, the possible ways to select variables is $d^{(\gamma_{1,\ell})-1}$. Denote those variables as $C^{j}_{i,k_{2}},\cdots,C^{j}_{i,k_{\gamma_{1,\ell}}}$.
    3. Step 3
      Pick up a specific $k_{\gamma_{1,\ell}+1}$ from $[p]$ except elements in $B_{j}$ and $B^{*}_{1,\ell}$. The possible ways to select $C^{j}_{ik_{(\gamma_{1,\ell})+1}}$ is $p-(|B_{j}|+|B^{*}_{1,\ell}|)$. Let $B^{*}_{2,\ell}$ denote the dependent index cell of $C^{j}_{ik_{(\gamma_{1,\ell})+1}}$.
    4. Step 4
      Set $\gamma_{2,\ell}$ as the number of variables included in the same dependent index cell as $C^{j}_{ik_{(\gamma_{1,\ell})+1}}$. As in Step 2, the possible ways to select variables is $d^{(\gamma_{2,\ell})-1}$. Denote those variables as $C^{j}_{i,k_{2}},\cdots,C^{j}_{i,k_{\gamma_{1,\ell}}}$.
    5. Step 5
      Iterate this process until we define $B^{*}_{\ell,\ell}$ and $\gamma_{\ell,\ell}$
    Because the possible ways to choose $\gamma_{1,\ell},\cdots,\gamma_{\ell,\ell}$ only depends on $q$, we have
    $N_{1}$ $\lesssim_{q}\sum_{\ell=1}^{\bar{\mathbf{q}}_{0}-\alpha-1}p^{\ell}d^{\sum_{\ell^{*}=1}^{\ell}(\gamma_{\ell^{*},\ell}-1)}$
    $=\sum_{\ell=1}^{\bar{\mathbf{q}}_{0}-\alpha-1}p^{\ell}d^{\bar{\mathbf{q}}_{0}-\ell}$
    where the last equality follows from the definition of $\gamma_{\ell^{*},\ell}$, i.e. $\sum_{\ell^{*}=1}^{\ell}\gamma_{\ell^{*},\ell}=\bar{\mathbf{q}}_{0}$. Moreover, because $d\leq p$ holds, for any $21\leq\ell\leq\bar{\mathbf{q}}_{0}-\alpha-$ 2, we have
    $p^{\ell}d^{\bar{\mathbf{q}}_{0}-\ell}\leq p^{\bar{\mathbf{q}}_{0}/2-1}d^{\bar{\mathbf{q}}_{0}/2+1}.$
    Therefore, we have
    $N_{1}\lesssim_{q}p^{\bar{\mathbf{q}}_{0}/2-1}d^{\bar{\mathbf{q}}_{0}/2+1}.$
    Apparently, the upper bound of the uniform expectations is $\max_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\in[p]^{\bar{\mathbf{q}}_{0}}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]$. Therefore, we have
    $A_{1}\lesssim_{q}p^{\bar{\mathbf{q}}_{0}/2-1}d^{\bar{\mathbf{q}}_{0}/2+1}\max_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\in[p]^{\bar{\mathbf{q}}_{0}}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right].$
  2. (ii)
    When there exists at least one dependent index cell that has only one element, that term becomes zero due to the expectation of that element. From the definition of $\ell$, there is no dependent index cell where more than two random variables belong. Thus, the expectation can be divided into pairs of random variables. Moreover, we only have ${}_{\bar{\mathbf{q}}_{0}}\rm C_{2}$ patters of pairs belonging to the same dependent cell from $C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}$. Hence, to derive the upper bound, all we have to do is to consider the case when $k_{t}$ and $k_{t+1}$ are in the same dependent index cell where $t$ is odd, i.e.
    $\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]$
    $=$ $\sum_{\begin{subarray}{c}\ell_{1}=1\\ \ell_{1}\neq j\end{subarray}}^{k}\sum_{\ell_{2}\neq\ell_{1},j}^{k}\cdots\sum_{\ell_{\alpha}\neq\ell_{1},\ell_{2},\cdots,\ell_{\alpha-1},j}^{k}\sum_{\{k_{1},k_{2}\}\in B_{\ell_{1}}}\cdots\sum_{\{k_{\bar{\mathbf{q}}_{0}-1},k_{\bar{\mathbf{q}}_{0}}\}\in B_{\ell_{\alpha}}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]$
    By definition of the dependnet structure, we have
    $\sum_{\begin{subarray}{c}\ell_{1}=1\\ \ell_{1}\neq j\end{subarray}}^{k}\sum_{\ell_{2}\neq\ell_{1},j}\cdots\sum_{\ell_{\alpha}\neq\ell_{1},\ell_{2},\cdots,\ell_{\alpha-1},j}\sum_{\{k_{1},k_{2}\}\in B_{\ell_{1}}}\cdots\sum_{\{k_{\bar{\mathbf{q}}_{0}-1},k_{\bar{\mathbf{q}}_{0}}\}\in B_{\ell_{\alpha}}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]$
    $=$ $\sum_{\begin{subarray}{c}\ell_{1}=1\\ \ell_{1}\neq j\end{subarray}}^{k}\sum_{\ell_{2}\neq\ell_{1},j}\cdots\sum_{\ell_{\alpha}\neq\ell_{1},\ell_{2},\cdots,\ell_{\alpha-1},j}\sum_{\{k_{1},k_{2}\}\in B_{\ell_{1}}}\cdots\sum_{\{k_{\bar{\mathbf{q}}_{0}-1},k_{\bar{\mathbf{q}}_{0}}\}\in B_{\ell_{\alpha}}}E\left[C^{j}_{ik_{1}}C^{j}_{ik_{2}}\right]\cdots E\left[C^{j}_{ik_{\bar{\mathbf{q}}_{0}-1}}C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]$
    $=$ $\left(\frac{1}{n}\right)^{\alpha}\sum_{\begin{subarray}{c}\ell_{1}=1\\ \ell_{1}\neq j\end{subarray}}^{k}\sum_{\ell_{2}\neq\ell_{1},j}\cdots\sum_{\ell_{\alpha}\neq\ell_{1},\ell_{2},\cdots,\ell_{\alpha-1},j}\bm{\iota}_{|B_{\ell_{1}}|}^{\top}\Sigma_{|B_{\ell_{1}}|}\bm{\iota}_{|B_{\ell_{1}}|}\bm{\iota}_{|B_{\ell_{2}}|}^{\top}\Sigma_{|B_{\ell_{2}}|}\bm{\iota}_{|B_{\ell_{2}}|}\cdots\bm{\iota}_{|B_{\ell_{\alpha}}|}^{\top}\Sigma_{|B_{\ell_{\alpha}}|}\bm{\iota}_{|B_{\ell_{\alpha}}|}$
    $\leq$ $\left(\frac{1}{n}\right)^{\alpha}(\max_{1\leq i\leq k}\lambda_{\max}(\Sigma_{|B_{i}|}))^{\alpha}\sum_{\begin{subarray}{c}\ell_{1}=1\\ \ell_{1}\neq j\end{subarray}}^{k}\sum_{\ell_{2}\neq\ell_{1},j}\cdots\sum_{\ell_{\alpha}\neq\ell_{1},\ell_{2},\cdots,\ell_{\alpha-1},j}|B_{\ell_{1}}|\cdots|B_{\ell_{\alpha}}|$
    $\leq$ $\left(\frac{p}{n}\right)^{\alpha}(\max_{1\leq i\leq k}\lambda_{\max}(\Sigma_{|B_{i}|}))^{\alpha}$
    where we define $\Sigma_{|B_{i}|}$ as a variance-covariance matrix of the cell $B_{i}$. The last inequality is valid from the definition of the partition. Let $\bm{\iota}_{|B_{i}|}$ denote a $|B_{i}|\times 1$ vector of ones. Therefore, we obtain
    $A_{2}\lesssim_{q,\tau}(\max_{1\leq i\leq k}\lambda_{\max}(\Sigma_{|B_{i}|}))^{\bar{\mathbf{q}}_{0}/2}.$
  3. (iii)
    In this case, at least one cell consists of just one element of $C^{j}_{i1},\cdots,C^{j}_{i\bar{\mathbf{q}}_{0}}$, which implies we have $E\left[C^{j}_{ik_{1}}\cdots C^{j}_{i\bar{\mathbf{q}}_{0}}\right]=0$. Therefore, $A_{3}$ is equal to 0.
In summary, the discussion above yields
$\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]$
$\lesssim_{q,\tau}(\max_{1\leq i\leq k}\lambda_{\max}(\Sigma_{|B_{i}|}))^{\bar{\mathbf{q}}_{0}/2}+p^{\bar{\mathbf{q}}_{0}/2-1}d^{\bar{\mathbf{q}}_{0}/2+1}\max_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\in[p]^{\bar{\mathbf{q}}_{0}}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right].$
As in the proof of Proposition 7, we have
$\max_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\in[p]^{\bar{\mathbf{q}}_{0}}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]\leq n^{-\bar{\mathbf{q}}_{0}/2}\max\{\max_{1\leq k\leq p}(\|X_{i,k}\|_{2^{(\bar{\mathbf{q}}_{0}/2)}})^{(\bar{\mathbf{q}}_{0})},\max_{1\leq k\leq p}(\|W_{i,k}\|_{2^{(\bar{\mathbf{q}}_{0}/2)}})^{(\bar{\mathbf{q}}_{0})}\}.$
Therefore, the required result holds. ∎